### v1

# Self-consistent constraints on the collision terms in quantum kinetic theory

Abstract
We derive self-consistent constraint conditions for collision terms in quantum kinetic theory within Wigner function formalism. We present simple specific solutions for the collision terms in quantum kinetic theory which is compatible with these constraints. We obtain the self-consistent quantum kinetic theory at relaxation-time approximation, with which we discuss spin polarization effects and electric charge separation effects in relativistic heavy-ion collisions.
1 Introduction
In 1872, Boltzmann, the founder of the kinetic theory, derived his famous classical kinetic equation from the physical point of view Boltzmann:1872 . The quantum analog of the Boltzmann equation was suggested by Uehling and Uhlenbeck in 1933 by purely phenomenological considerationUehling:1933 . It was not until 1946 that Bogolyubov, Born, Green, Kirkwood Yvon and others derived the kinetic equations within the mathematical formalism from the Liouville equation Bogolyubov:1946 . The quantum kinetic equations based on quantum field theory was derived by Martin and Schwinger Martin:1959jp ; Schwinger:1960qe , Kadanoff and Baym Baym:1961zz , Keldysh Keldysh:1964ud with the method of non-equilibrium Green functions. In the 1980s and 1990s , the gauge invariant quantum kinetic theories based on QED or QCD were derived Heinz:1983nx ; Heinz:1985vf ; Elze:1986qd ; Elze:1986hq ; Vasak:1987um ; Zhuang:1995pd ; Ochs:1998qj ; Zhuang:1998bqx ; Wang:2001dm ; Wang:2002qe in order to describe non-equilibrium and quantum effects in upcoming relativistic heavy-ion collisions at that time. In the 2000s, the pioneering work Liang:2004ph ; Liang:2004xn ; Kharzeev:2007jp ; Fukushima:2008xe ; Kharzeev:2007tn ; Gao:2007bc ; Becattini:2007sr initiated the research on the chiral and spin effects in relativistic heavy-ion collisions, which lead to further development on the derivation of quntum kinetic theory with chiral or spin degree of freedom Gao:2012ix ; Stephanov:2012ki ; Son:2012zy ; Chen:2012ca ; Manuel:2013zaa ; Manuel:2014dza ; Chen:2014cla ; Chen:2015gta ; Hidaka:2016yjf ; Mueller:2017lzw ; Huang:2018wdl ; Gao:2018wmr ; Liu:2018xip ; Gao:2019znl ; Weickgenannt:2019dks ; Hattori:2019ahi ; Wang:2019moi ; Lin:2019fqo ; Li:2019qkf ; Sheng:2020oqs ; Guo:2020zpa ; Weickgenannt:2020aaf ; Hayata:2020sqz ; Weickgenannt:2021cuo ; Sheng:2021kfc ; Fauth:2021nwe ; Luo:2021uog ; ChenShiLe:2021guf ; Lin:2021mvw ; Chen:2021azy ; Sheng:2022ssd ; Fang:2022ttm ; Wagner:2022amr ; Ma:2022ins ; Wagner:2023cct .
Collision terms in kinetic equation play a central role in describing the evolution from non-equilibrium to equilibrium, and can be derived from the BBGKY equation or hierarchyBogolyubov:1946 by introducing certain additional assumptions. As we all know, in general, the collision terms should satisfy the principle of detailed balance and the constraints of charge conservation or energy-momentum conservation. In this paper, we will demonstrate that the collision terms in quantum kinetic theory must satisfy some extra constraints totally determined by the self-consistent nature of the quantum kinetic equation. With this self-consistent constraints, we will give non-trivial specific solutions for the collision terms in quantum kinetic theory. We put the specific solution in the form of relaxation-time approximation and use it to discuss spin polarization and electric charge separation in relativistic heavy-ion collisions.
We will use the convention for the metric $g_{\mu\nu}=$ diag(1,1,1,1)$1$(1,-1,-1,-1)( 1 , - 1 , - 1 , - 1 ), Levi Civita tensor $\epsilon^{0123}=1$. We choose natural units such that $\hbar=c=1$ unless otherwise stated.
2 Wigner function formalism
The quantum kinetic theory can be built with different formalisms. In this work, we will adopt the Wigner function formalism Heinz:1983nx ; Heinz:1985vf ; Elze:1986qd ; Elze:1986hq ; Vasak:1987um ; Zhuang:1995pd ; Ochs:1998qj ; Zhuang:1998bqx ; Wang:2001dm ; Wang:2002qe and restrict ourselves to the system controlled by quantum electrodynamics. In quantum electrodynamics, we define gauge invariant density operator Elze:1986qd as the following,
$\rho_{ab}\left(x,y\right)=\bar{\psi}_{b}(x)e^{\frac{y}{2}\cdot D^{\dagger}}e^{-\frac{y}{2}\cdot D}\psi_{a}(x),$ (1)
where $\psi_{a}$ and $\bar{\psi}_{b}$ represent the electron’s spinor field with the spinor indices $a$ and $b$ running from 1 to 4. The covariant derivative $D$ or conjugate $D^{\dagger}$ in the covariant translation operator are given by
$D_{\mu}=\partial_{\mu}^{x}+iA_{\mu}(x),\ \ \ D_{\mu}^{\dagger}=\partial^{x\dagger}_{\mu}-iA_{\mu}(x),$ (2)
where we have absorbed the electric charge $e$ into the gauge potential $A_{\mu}$, and the derivatives, $\partial_{\mu}^{x}$ and $\partial^{x\dagger}_{\mu}$, denote acting on the right and the left with respect to the coordinate $x$ , respectively. We note that the density operator $\rho$ in Hilbert space is valued as a $4\times 4$ matrix in spinor space.
The Wigner function is defined as the Fourier transformation of the ensemble averaging of the density operator by
$W(x,p)=\int\frac{d^{4}y}{(2\pi)^{4}}e^{-ip\cdot y}\langle\rho\left(x,y\right)\rangle,$ (3)
where the brackets denote the ensemble average. After ensemble average, the Wigner function is only a $4\times 4$ matrix in spinor space. The Wigner equation satisfied by $W(x,p)$ can be derived from the Dirac equation Elze:1986qd and put in the following form
$\left[m-\gamma_{\mu}\left(p^{\mu}+\frac{1}{2}i\partial^{\mu}_{x}\right)\right]W(x,p)$ $=$ $\gamma_{\mu}\left[\frac{1}{2}i{C}^{\mu}(x,p)+\Delta{C}^{\mu}(x,p)\right],$ (4)
where $C^{\mu}(x,p)$ and $\Delta C^{\mu}(x,p)$ are both four-vectors in Minkowski space but also $4\times 4$ matrices in spinor space with the element defined as
${C}^{\mu}_{ab}$ $=$ $j_{0}\left(\frac{\Delta}{2}\right)\partial_{\nu}^{p}\int\frac{d^{4}y}{(2\pi)^{4}}e^{-ip\cdot y}\langle\bar{\psi}_{b}(x)e^{\frac{y}{2}\cdot D^{\dagger}}F^{\nu\mu}\left(x\right)e^{-\frac{y}{2}\cdot D}\psi_{a}(x)\rangle,$ (5)
$\Delta{C}^{\mu}_{ab}$ $=$ $\frac{1}{2}j_{1}\left(\frac{\Delta}{2}\right)\partial_{\nu}^{p}\int\frac{d^{4}y}{(2\pi)^{4}}e^{-ip\cdot y}\langle\bar{\psi}_{b}(x)e^{\frac{y}{2}\cdot D^{\dagger}}F^{\nu\mu}\left(x\right)e^{-\frac{y}{2}\cdot D}\psi_{a}(x)\rangle,$ (6)
where $j_{0}$ and $j_{1}$ are zeroth- and first-order spherical Bessel functions, respectively, and the triangle operator $\Delta\equiv\partial_{p}\cdot\partial_{x}$ denotes the mixed derivative. The electromagnetic field tensor $F_{\nu\mu}\left(x\right)$ is defined as usual
$F_{\mu\nu}(x)=\partial_{\mu}^{x}A_{\nu}(x)-\partial_{\nu}^{x}A_{\mu}(x).$ (7)
It should be noted that the derivative $\partial_{x}$ with respect to $x$ in triangle operator $\Delta$ only acts on the electromagnetic field tensor.
We can regard the Wigner function $W(x,p)$ as one-body function while $C^{\mu}(x,p)$ and $\Delta C^{\mu}(x,p)$ as two-body functions. Thus the Wigner equation (4) shows that the equation of one-body function depends on the two-body function, which is so-called the BBGKY hierarchyBogolyubov:1946 . Within the mean field approximation, we can pull the tensor $F_{\nu\mu}\left(x\right)$ out of the ensemble average and the BBGKY hierarchy truncates at the one-body function $W(x,p)$ with the result of the quantum Vlasov equation. For the general quantum fields, the function $C^{\mu}(x,p)$ and $\Delta C^{\mu}(x,p)$ will lead to collision terms, which are the main subject of our present work as indicated from the title of this paper. However, it should be emphasized that $C^{\mu}(x,p)$ and $\Delta C^{\mu}(x,p)$ include not only collisions terms but also all possible terms, such as mean field terms and so on. Hence, in our present work, the collision terms denote the full terms $C^{\mu}(x,p)$ and $\Delta C^{\mu}(x,p)$. We will demonstrate that this general collision terms must satisfy some self-consistent constraints determined by the Wigner equation itself. From now on, we will simply name $C^{\mu}(x,p)$ and $\Delta C^{\mu}(x,p)$ or the associated function as the collision function.
The Wigner function can be expanded in the 16 covariant matrices $1$, $i\gamma^{5}$, $\gamma_{\nu}$, $\gamma^{5}\gamma_{\nu}$, and $\sigma_{\mu\nu}=i[\gamma_{\mu},\gamma_{\nu}]/2$, i.e.,
$W$ $=$ $\frac{1}{4}\left[\mathscr{F}+i\gamma^{5}\mathscr{P}+\gamma_{\nu}\mathscr{V}^{\nu}+\gamma^{5}\gamma_{\nu}\mathscr{A}^{\nu}+\frac{1}{2}\sigma_{\mu\nu}\mathscr{S}^{\mu\nu}\right],$ (8)
with the real coefficients representing scalar $\mathscr{F}$, pseudoscalar $\mathscr{P}$, vector $\mathscr{V}^{\nu}$, axial vector $\mathscr{A}^{\nu}$ and antisymmetric tensor $\mathscr{S}^{\mu\nu}$ components, respectively. In general, the quantum kinetic theory for the fermions with arbitrary mass might exhibit very different forms if we choose different Wigner functions as independent distribution functionsGao:2019znl ; Weickgenannt:2019dks ; Hattori:2019ahi ; Wang:2019moi ; Ma:2022ins . In this work, we will follow the formalism of generalized chiral kinetic theory (GCKT) derived in Ref.Ma:2022ins . The GCKT can be reduced into the chiral kinetic theory with a smooth transition from massive to massless fermions. In GCKT, we need introduce chiral Wigner function via
$\mathscr{J}^{\nu}_{s}=\frac{1}{2}\left(\mathscr{V}^{\nu}+s\mathscr{A}^{\nu}\right),$ (9)
where $s=+1/-1$ denotes the chirality with right-handed/left-handed component. Similarly, we can expand the collision functions as
${C}^{\mu}$ $=$ $\frac{1}{4}\left[\mathscr{C}^{\mu}+i\gamma^{5}\mathscr{C}^{\mu}_{5}+\gamma_{\nu}\mathscr{C}^{\mu\nu}+\gamma^{5}\gamma_{\nu}\mathscr{C}^{\mu\nu}_{5}+\frac{1}{2}\sigma_{\alpha\beta}\mathscr{C}^{\mu\alpha\beta}\right],$ (10)
$\Delta{C}^{\mu}$ $=$ $\frac{1}{4}\left[\Delta\mathscr{C}^{\mu}+i\gamma^{5}\Delta\mathscr{C}^{\mu}_{5}+\gamma_{\nu}\Delta\mathscr{C}^{\mu\nu}+\gamma^{5}\gamma_{\nu}\Delta\mathscr{C}^{\mu\nu}_{5}+\frac{1}{2}\sigma_{\alpha\beta}\Delta\mathscr{C}^{\mu\alpha\beta}\right],$ (11)
with corresponding chiral collision functions
$\ \ \mathscr{C}^{\mu\nu}_{s}=\frac{1}{2}\left(\mathscr{C}^{\mu\nu}+s\mathscr{C}^{\mu\nu}_{5}\right),\ \ \Delta\mathscr{C}^{\mu\nu}_{s}=\frac{1}{2}\left(\Delta\mathscr{C}^{\mu\nu}+s\Delta\mathscr{C}^{\mu\nu}_{5}\right).$ (12)
With these decomposition and chiral functions, the Wigner equation can be cast into
$m\mathscr{F}$ $=$ $2p_{\mu}\mathscr{J}^{\mu}_{s}+\hbar\Delta\mathscr{C}^{\mu}_{s,\mu},$ (13)
$-sm\mathscr{P}$ $=$ $\hbar\left(\partial_{\mu}^{x}\mathscr{J}^{\mu}_{s}+\mathscr{C}^{\mu}_{s,\mu}\right),$ (14)
$\frac{1}{2}m\epsilon_{\mu\nu\alpha\beta}\mathscr{S}^{\alpha\beta}$ $=$ $\hbar\epsilon_{\mu\nu\alpha\beta}\left(\partial^{\alpha}_{x}\mathscr{J}^{\beta}_{s}+\mathscr{C}^{\alpha\beta}_{s}\right)$ (15)
$+s\left[2(p_{\mu}\mathscr{J}_{s,\nu}-p_{\nu}\mathscr{J}_{s,\mu})+\hbar\left(\Delta\mathscr{C}_{s,\mu\nu}-\Delta\mathscr{C}_{s,\nu\mu}\right)\right],$
$2m\sum_{s}\mathscr{J}_{s,\mu}$ $=$ $2p_{\mu}\mathscr{F}+\hbar\Delta\mathscr{C}_{\mu}+\hbar\left(\partial^{\nu}_{x}\mathscr{S}_{\mu\nu}+\mathscr{C}^{\nu}_{\ \mu\nu}\right),$ (16)
$-2m\sum_{s}s\mathscr{J}_{s,\mu}$ $=$ $\epsilon_{\mu\nu\alpha\beta}\left(p^{\nu}\mathscr{S}^{\alpha\beta}+\frac{\hbar}{2}\Delta\mathscr{C}^{\nu\alpha\beta}\right)-\hbar\left(\partial_{\mu}^{x}\mathscr{P}+\mathscr{C}_{5,\mu}\right),$ (17)
00 0 $=$ $\hbar\left(\partial_{\mu}^{x}\mathscr{F}+\mathscr{C}_{\mu}\right)-2p^{\nu}\mathscr{S}_{\mu\nu}-\hbar\Delta\mathscr{C}^{\nu}_{\ \mu\nu},$ (18)
00 0 $=$ $\frac{1}{2}\hbar\epsilon_{\mu\nu\alpha\beta}\left(\partial^{\nu}_{x}\mathscr{S}^{\alpha\beta}+\mathscr{C}^{\nu\alpha\beta}\right)+2p_{\mu}\mathscr{P}+\hbar\Delta\mathscr{C}_{5,\mu},$ (19)
where we have recovered the $\hbar$ dependence so that we can make semiclassical expansion in the next section.
3 Self-consistent constraints on collision terms
The Wigner equations given in Eqs.(13)-(19) are very complicated and totally coupled with each other. It has been verified in Gao:2018wmr ; Gao:2019znl ; Ma:2022ins that these equations under mean field approximation can be reduced to very simple form, in which only very few Wigner functions and equations are independent. Some other Wigner functions can be derivative directly from the independent functions we chosen and some Wigner equations are fulfilled automatically and redundant. In this section, we will generalize this disentangling method in mean field approximation to general quantum field here. In order to achieve this goal, we resort to the semiclassical $\hbar$ expansion. Besides the explicit $\hbar$ dependence shown in Eqs.(13)-(19), the collision functions also have explicit $\hbar$ expansion when we recover the $\hbar$ dependence after we replace $\Delta$ by $\hbar\Delta$. In current work, we will restrict ourselves to the Wigner equation up to the first order. Then we only need to keep the leading contribution from $\Delta$ expansion:
${C}^{\mu}_{ab}$ $=$ $\partial_{\nu}^{p}\int\frac{d^{4}y}{(2\pi)^{4}}e^{-ip\cdot y}\langle\bar{\psi}_{b}(x)e^{\frac{y}{2}\cdot D^{\dagger}}F^{\nu\mu}\left(x\right)e^{-\frac{y}{2}\cdot D}\psi_{a}(x)\rangle,$ (20)
$\Delta{C}^{\mu}_{ab}$ $=$ $\frac{\hbar}{6}\partial_{\lambda}^{p}\partial_{\nu}^{p}\int\frac{d^{4}y}{(2\pi)^{4}}e^{-ip\cdot y}\langle\bar{\psi}_{b}(x)e^{\frac{y}{2}\cdot D^{\dagger}}\partial_{\lambda}F^{\nu\mu}\left(x\right)e^{-\frac{y}{2}\cdot D}\psi_{a}(x)\rangle.$ (21)
According to the decomposition in Eq.(10), we have the expressions up to this order,
${\mathscr{C}^{\nu}}$$\equiv$ $\partial_{\lambda}^{p}\int\frac{d^{4}y}{(2\pi)^{4}}e^{-ip\cdot y}\left\langle\bar{\psi}e^{\frac{y}{2}\cdot D^{\dagger}}F^{\nu\lambda}e^{-\frac{y}{2}\cdot D}\psi\right\rangle,$ (22)
${\mathscr{C}^{\nu}_{5}}$$\equiv$ $-i\partial_{\lambda}^{p}\int\frac{d^{4}y}{(2\pi)^{4}}e^{-ip\cdot y}\left\langle\bar{\psi}e^{\frac{y}{2}\cdot D^{\dagger}}F^{\nu\lambda}\gamma^{5}e^{-\frac{y}{2}\cdot D}\psi\right\rangle,$ (23)
$\mathscr{C}^{\mu\nu}_{s}$$\equiv$ $\frac{1}{2}\partial_{\lambda}^{p}\int\frac{d^{4}y}{(2\pi)^{4}}e^{-ip\cdot y}\left\langle\bar{\psi}e^{\frac{y}{2}\cdot D^{\dagger}}F^{\mu\lambda}\gamma^{\nu}(1+s\gamma^{5})e^{-\frac{y}{2}\cdot D}\psi\right\rangle,$ (24)
${\mathscr{C}^{\nu\alpha\beta}}$$\equiv$ $\partial_{\lambda}^{p}\int\frac{d^{4}y}{(2\pi)^{4}}e^{-ip\cdot y}\left\langle\bar{\psi}e^{\frac{y}{2}\cdot D^{\dagger}}F^{\nu\lambda}\sigma^{\alpha\beta}e^{-\frac{y}{2}\cdot D}\psi\right\rangle,$ (25)
$\Delta{\mathscr{C}^{\nu}}$$\equiv$ $\frac{\hbar}{6}\partial_{\lambda}^{p}\partial_{\kappa}^{p}\int\frac{d^{4}y}{(2\pi)^{4}}e^{-ip\cdot y}\left\langle\bar{\psi}e^{\frac{y}{2}\cdot D^{\dagger}}\partial^{\kappa}_{x}F^{\nu\lambda}e^{-\frac{y}{2}\cdot D}\psi\right\rangle,$ (26)
$\Delta{\mathscr{C}^{\nu}_{5}}$$\equiv$ $-\frac{i\hbar}{6}\partial_{\lambda}^{p}\partial_{\kappa}^{p}\int\frac{d^{4}y}{(2\pi)^{4}}e^{-ip\cdot y}\left\langle\bar{\psi}e^{\frac{y}{2}\cdot D^{\dagger}}\partial^{\kappa}_{x}F^{\nu\lambda}\gamma^{5}e^{-\frac{y}{2}\cdot D}\psi\right\rangle,$ (27)
$\Delta\mathscr{C}^{\mu\nu}_{s}$$\equiv$ $\frac{\hbar}{12}\partial_{\lambda}^{p}\partial_{\kappa}^{p}\int\frac{d^{4}y}{(2\pi)^{4}}e^{-ip\cdot y}\left\langle\bar{\psi}e^{\frac{y}{2}\cdot D^{\dagger}}\partial^{\kappa}_{x}F^{\mu\lambda}\gamma^{\nu}(1+s\gamma^{5})e^{-\frac{y}{2}\cdot D}\psi\right\rangle,$ (28)
$\Delta{\mathscr{C}^{\nu\alpha\beta}}$$\equiv$ $\frac{\hbar}{6}\partial_{\lambda}^{p}\partial_{\kappa}^{p}\int\frac{d^{4}y}{(2\pi)^{4}}e^{-ip\cdot y}\left\langle\bar{\psi}e^{\frac{y}{2}\cdot D^{\dagger}}\partial^{\kappa}_{x}F^{\nu\lambda}\sigma^{\alpha\beta}e^{-\frac{y}{2}\cdot D}\psi\right\rangle.$ (29)
where we have suppressed the arguments $x$ of the fields $\bar{\psi}$, $\psi$, and $F^{\nu\lambda}$ for simplicity of notations. We should note that even only leading term defined above can also contribute at any higher order which is implicit in the ensemble average of the operators, i.e.,
$\mathscr{C}^{\mu}_{ab}=\sum_{k=0}^{\infty}\hbar^{k}\mathscr{C}^{(k)\mu}_{ab},\ \ \ \Delta\mathscr{C}^{\mu}_{ab}=\sum_{k=0}^{\infty}\hbar^{k+1}\Delta\mathscr{C}^{(k)\mu}_{ab},$ (30)
just as we do for the Wigner function
$W(x,p)=\sum_{k=0}^{\infty}\hbar^{k}W^{(k)}(x,p).$ (31)
In order to disentangle the Wigner equations further, we introduce time-like 4-vector $n_{\mu}$ with normalization condition $n^{2}=1$. In this work, we will assume that $n_{\mu}$ can be a function of coordinates $x^{\mu}$. Then we can decompose any 4-vector $X^{\mu}$ as
$X^{\mu}=X_{n}n^{\mu}+\bar{X}^{\mu},$ (32)
where $X_{n}=X\cdot n$ and $\bar{X}^{\mu}=\Delta^{\mu\nu}X_{\nu}$ with $\Delta^{\mu\nu}=g^{\mu\nu}-n^{\mu}n^{\nu}$. We can also decompose the antisymmetric tensor as
$\mathscr{S}^{\mu\nu}$ $=$ $\mathscr{K}^{\mu}n^{\nu}-\mathscr{K}^{\nu}n^{\mu}+\epsilon^{\mu\nu\rho\sigma}n_{\rho}\mathscr{M}_{\sigma},$ (33)
with the evident relations
$\mathscr{K}^{\mu}$ $=$ $\mathscr{S}^{\mu\nu}n_{\nu},\ \ \mathscr{M}^{\mu}=\frac{1}{2}\epsilon^{\mu\nu\rho\sigma}n_{\nu}\mathscr{S}_{\rho\sigma}.$ (34)
It is also convenient to define the totally space-like antisymmetric tensor as
$\bar{\epsilon}_{\mu\alpha\beta}=\epsilon_{\mu\nu\alpha\beta}n^{\nu}.$ (35)
3.1 Zeroth-order result
At zeroth order, all the collision functions vanish in the Wigner equations Eqs.(13)-(19). After the time-like and space-like decomposition according to the time-like vector $n^{\mu}$ for the Wigner functions and equations, the zeroth-order result can be presented as the following
$m{\mathscr{F}}^{(0)}$ $=$ $2\left(p_{n}{\mathscr{J}}^{(0)}_{s,n}+\bar{p}_{\mu}\bar{\mathscr{J}}^{(0)\mu}_{s}\right),$ (36)
$-sm{\mathscr{P}}^{(0)}$ $=$ $0,$ (37)
$m{\mathscr{M}}^{(0)}_{\mu}$ $=$ $2s\left(\bar{p}_{\mu}{\mathscr{J}}_{s,n}^{(0)}-p_{n}\bar{\mathscr{J}}_{s,\mu}^{(0)}\right),$ (38)
$m\bar{\epsilon}_{\mu\nu\alpha}{\mathscr{K}}^{(0)\alpha}$ $=$ $2s\left(\bar{p}_{\mu}\bar{\mathscr{J}}_{s,\nu}^{(0)}-\bar{p}_{\nu}\bar{\mathscr{J}}_{s,\mu}^{(0)}\right),$ (39)
$2m\sum_{s}{\mathscr{J}}_{s,n}^{(0)}$ $=$ $2p_{n}{\mathscr{F}}^{(0)},$ (40)
$2m\sum_{s}\bar{\mathscr{J}}_{s,\mu}^{(0)}$ $=$ $2\bar{p}_{\mu}{\mathscr{F}}^{(0)},$ (41)
$-2m\sum_{s}s{\mathscr{J}}_{s,n}^{(0)}$ $=$ $-2\bar{p}^{\mu}{\mathscr{M}}^{(0)}_{\mu},$ (42)
$-2m\sum_{s}s\bar{\mathscr{J}}_{s,\mu}^{(0)}$ $=$ $2\left(p_{n}\mathscr{M}^{(0)}_{\mu}+\bar{\epsilon}_{\mu\nu\alpha}\bar{p}^{\nu}\mathscr{K}^{(0)\alpha}\right),$ (43)
00 0 $=$ $-2p^{\nu}{\mathscr{K}}_{\nu}^{(0)},$ (44)
00 0 $=$ $-2\left(p_{n}\mathscr{K}^{(0)}_{\mu}-\bar{\epsilon}_{\mu\nu\alpha}\bar{p}^{\nu}\mathscr{M}^{(0)\alpha}\right),$ (45)
00 0 $=$ $2p_{n}{\mathscr{P}}^{(0)},$ (46)
00 0 $=$ $2\bar{p}_{\mu}{\mathscr{P}}^{(0)}.$ (47)
The sequence of the equations listed above is in correspondence with the one in Eqs.(13)-(19). The equation (38) is from the tensor equation (15) with the time-like component $\nu$ and space-like component $\mu$ while the equation (39) is from the space-like and space-like components for both $\mu$ and $\nu$. The equations (40)/(41), (42)/(43), (44)/(45), and (46)/(47) are from the time-like/space-like components of Eqs.(16),(17), (18), and (19), respectively
We will choose ${\mathscr{J}}_{s,n}$ and $\mathscr{M}_{\mu}$ as the independent Wigner functions and the other Wigner functions as the derived functions. From the Eqs.(38), (40),(45) and (46), we can express the Wigner functions $\bar{\mathscr{J}}_{s,\mu}^{(0)}$, ${\mathscr{F}}^{(0)}$, $\mathscr{K}^{(0)}_{\mu}$ and ${\mathscr{P}}^{(0)}$ in terms of independent functions ${\mathscr{J}}_{s,n}^{(0)}$ and ${\mathscr{M}}^{(0)}_{\mu}$, respectively,
$\bar{\mathscr{J}}_{s,\mu}^{(0)}$ $=$ $\frac{\bar{p}_{\mu}}{p_{n}}{\mathscr{J}}_{s,n}^{(0)}-\frac{sm}{2p_{n}}{\mathscr{M}}^{(0)}_{\mu},$ (48)
${\mathscr{F}}^{(0)}$ $=$ $\frac{m}{p_{n}}\sum_{s}{\mathscr{J}}_{s,n}^{(0)},$ (49)
$\mathscr{K}^{(0)\mu}$ $=$ $\frac{1}{p_{n}}\bar{\epsilon}^{\mu\nu\alpha}\bar{p}_{\nu}\mathscr{M}_{\alpha}^{(0)},$ (50)
${\mathscr{P}}^{(0)}$ $=$ $0$ (51)
Substituting Eq.(50) into Eq.(33), we can obtain the antisymmetric Wigner function
$\mathscr{S}^{(0)\mu\nu}$ $=$ $\frac{1}{p_{n}}\epsilon^{\mu\nu\alpha\beta}p_{\alpha}\mathscr{M}_{\beta}^{(0)}.$ (52)
The Eq.(42) gives the constraint condition for ${\mathscr{M}}^{(0)}_{\mu}$
$\bar{p}^{\mu}{\mathscr{M}}^{(0)}_{\mu}$ $=$ $m\sum_{s}s{\mathscr{J}}_{s,n}^{(0)}.$ (53)
This constraint and the one $n^{\mu}{\mathscr{M}}^{(0)}_{\mu}=0$ directly from the definition (34) means only two components are independent for ${\mathscr{M}}^{(0)}_{\mu}$. It is convenient to decompose $\mathscr{M}^{(0)}_{\mu}$ into the longitudinal and transverse parts with respect to the momentum $\bar{p}^{\mu}$
$\mathscr{M}^{(0)}_{\mu}=\mathscr{M}^{(0)}_{\parallel\mu}+\mathscr{M}^{(0)}_{\perp\mu},\ \ \ \textrm{with}\ \ \ \mathscr{M}^{(0)}_{\parallel\mu}=\frac{m\bar{p}_{\mu}}{\bar{p}^{2}}\sum_{s}s{\mathscr{J}}_{s,n}^{(0)}\ \ \textrm{and}\ \ \bar{p}^{\mu}\mathscr{M}^{(0)}_{\perp\mu}=0.$ (54)
We note that the longitudinal part $\mathscr{M}^{\mu}_{\parallel}$ along the direction of $\bar{p}^{\mu}$ is totally determined by the function ${\mathscr{J}}_{s,n}^{(0)}$ and the independent part is only transverse component $\mathscr{M}^{\mu}_{\perp}$.
Substituting the expressions given in (48)-(51) into Eqs.(36) and (43) and using the constraint (53) , we obtain the on-shell conditions for ${\mathscr{J}}_{s,n}^{(0)}$ and ${\mathscr{M}}^{(0)}_{\mu}$, respectively,
$\left(p^{2}-m^{2}\right)\frac{{\mathscr{J}}_{s,n}^{(0)}}{p_{n}}$ $=$ $0,$ (55)
$\left(p^{2}-m^{2}\right)\frac{{\mathscr{M}}_{\mu}^{(0)}}{p_{n}}$ $=$ $0$ (56)
The expressions take the general form of
${\mathscr{J}}_{s,n}^{(0)}$ $=$ $p_{n}{\mathcal{J}}_{s,n}^{(0)}\delta(p^{2}-m^{2}),$ (57)
${\mathscr{M}}_{\mu}^{(0)}$ $=$ $p_{n}{\mathcal{M}}_{\mu}^{(0)}\delta(p^{2}-m^{2}),$ (58)
where ${\mathcal{J}}_{s,n}^{(0)}$ and ${\mathcal{M}}_{\mu}^{(0)}$ are both regular functions of $x^{\mu}$ and $p^{\mu}$ at $p^{2}-m^{2}=0$.
Once the equations (48)-(51) hold , it is trivial to verify that the Eqs.(37), (39), (41), (44) and (47) are all satisfied automatically.
At zeroth order, we have no either constraints for the collision function or the kinetic equation for independent Wigner functions ${\mathscr{J}}_{s,n}^{(0)}$ and ${\mathscr{M}}^{(0)}_{\mu}$.
3.2 First-order result
The collision terms and kinetic equation begin to appear at first order. Similar to the procedure we carried out at zeroth order, the first-order Wigner equations can be obtained from Eqs.(13)-(19)
$m{\mathscr{F}}^{(1)}$ $=$ $2\left(p_{n}{\mathscr{J}}^{(1)}_{s,n}+\bar{p}_{\mu}\bar{\mathscr{J}}^{(1)\mu}_{s}\right),$ (59)
$-sm{\mathscr{P}}^{(1)}$ $=$ $\partial^{x}_{\mu}{\mathscr{J}}^{(0)\mu}_{s}+\mathscr{C}^{(0)\mu}_{s,\mu},$ (60)
$m{\mathscr{M}}^{(1)}_{\mu}$ $=$ $2s\left(\bar{p}_{\mu}{\mathscr{J}}_{s,n}^{(1)}-p_{n}\bar{\mathscr{J}}_{s,\mu}^{(1)}\right)+\bar{\epsilon}_{\mu\alpha\beta}\partial_{x}^{\alpha}{\mathscr{J}}^{(0)\beta}_{s}+\bar{\epsilon}_{\mu\alpha\beta}\mathscr{C}^{(0)\alpha\beta}_{s},$ (61)
$m\bar{\epsilon}_{\mu\nu\alpha}{\mathscr{K}}^{(1)\alpha}$ $=$ $2s\left(\bar{p}_{\mu}\bar{\mathscr{J}}_{s,\nu}^{(1)}-\bar{p}_{\nu}\bar{\mathscr{J}}_{s,\mu}^{(1)}\right)+\bar{\epsilon}_{\mu\nu\alpha}n_{\beta}\left(\partial_{x}^{\alpha}{\mathscr{J}}^{(0)\beta}_{s}-\partial_{x}^{\beta}{\mathscr{J}}^{(0)\alpha}_{s}\right)$ (62)
$+\bar{\epsilon}_{\mu\nu\alpha}n_{\beta}\left(\mathscr{C}^{(0)\alpha\beta}_{s}-\mathscr{C}^{(0)\beta\alpha}_{s}\right),$
$2m\sum_{s}{\mathscr{J}}_{s,n}^{(1)}$ $=$ $2p_{n}{\mathscr{F}}^{(1)}+n^{\mu}\partial_{x}^{\nu}{\mathscr{S}}_{\mu\nu}^{(0)}+n^{\mu}{\mathscr{C}^{(0)\nu}}_{\mu\nu},$ (63)
$2m\sum_{s}\bar{\mathscr{J}}_{s,\mu}^{(1)}$ $=$ $2\bar{p}_{\mu}{\mathscr{F}}^{(1)}+\Delta_{\mu}^{\lambda}\partial_{x}^{\nu}{\mathscr{S}}_{\lambda\nu}^{(0)}+\Delta_{\mu}^{\lambda}{\mathscr{C}^{(0)\nu}}_{\lambda\nu},$ (64)
$-2m\sum_{s}s{\mathscr{J}}_{s,n}^{(1)}$ $=$ $-2\bar{p}_{\mu}{\mathscr{M}}^{(1)\mu}-n^{\mu}\partial^{x}_{\mu}{\mathscr{P}}^{(0)}-\mathscr{C}^{(0)}_{5,n},$ (65)
$-2m\sum_{s}s\bar{\mathscr{J}}_{s,\mu}^{(1)}$ $=$ $2\left(p_{n}\mathscr{M}^{(1)}_{\mu}+\bar{\epsilon}_{\mu\nu\alpha}\bar{p}^{\nu}\mathscr{K}^{(1)\alpha}\right)-\Delta_{\mu}^{\lambda}\partial^{x}_{\lambda}{\mathscr{P}}^{(0)}-\bar{\mathscr{C}^{(0)}}_{5,\mu},$ (66)
00 0 $=$ $2p^{\nu}{\mathscr{K}}_{\nu}^{(1)}+n^{\mu}\partial_{\mu}^{x}{\mathscr{F}}^{(0)}+\mathscr{C}^{(0)}_{n},$ (67)
00 0 $=$ $-2\left(p_{n}\mathscr{K}^{(1)}_{\mu}-\bar{\epsilon}_{\mu\nu\alpha}\bar{p}^{\nu}\mathscr{M}^{(1)\alpha}\right)+\Delta_{\mu}^{\lambda}\partial_{\lambda}^{x}{\mathscr{F}}^{(0)}+\bar{\mathscr{C}}^{(0)}_{\mu},$ (68)
00 0 $=$ $2p_{n}{\mathscr{P}}^{(1)}-\frac{1}{2}\bar{\epsilon}_{\nu\alpha\beta}\partial^{\nu}_{x}{\mathscr{S}}^{(0)\alpha\beta}-\frac{1}{2}\bar{\epsilon}_{\nu\alpha\beta}\mathscr{C}^{(0)\nu\alpha\beta},$ (69)
00 0 $=$ $2\bar{p}_{\mu}{\mathscr{P}}^{(1)}+\frac{1}{2}\Delta_{\mu}^{\lambda}\epsilon_{\lambda\nu\alpha\beta}\partial^{\nu}_{x}{\mathscr{S}}^{(0)\alpha\beta}+\frac{1}{2}\Delta_{\mu}^{\lambda}\epsilon_{\lambda\nu\alpha\beta}\mathscr{C}^{(0)\nu\alpha\beta}.$ (70)
From the Eqs.(61), (63),(68), and (69), we can express the functions $\bar{\mathscr{J}}_{s,\mu}^{(1)}$, ${\mathscr{F}}^{(1)}$, $\mathscr{K}^{(1)}_{\mu}$ and ${\mathscr{P}}^{(1)}$ in terms of independent functions ${\mathscr{J}}_{s,n}^{(1)}$ and ${\mathscr{M}}^{(1)}_{\mu}$, respectively,
$\bar{\mathscr{J}}_{s,\mu}^{(1)}$ $=$ $\frac{\bar{p}_{\mu}}{p_{n}}{\mathscr{J}}_{s,n}^{(1)}-\frac{sm}{2p_{n}}{\mathscr{M}}^{(1)}_{\mu}+\frac{s}{2p_{n}}\bar{\epsilon}_{\mu\alpha\beta}\partial_{x}^{\alpha}{\mathscr{J}}^{(0)\beta}_{s}+\frac{s}{2p_{n}}\bar{\epsilon}_{\mu\alpha\beta}\mathscr{C}^{(0)\alpha\beta}_{s},$ (71)
${\mathscr{F}}^{(1)}$ $=$ $\frac{m}{p_{n}}\sum_{s}{\mathscr{J}}_{s,n}^{(1)}-\frac{1}{2p_{n}}n^{\mu}\partial_{x}^{\nu}{\mathscr{S}}_{\mu\nu}^{(0)}-\frac{1}{2p_{n}}n^{\mu}{\mathscr{C}^{(0)\nu}}_{\mu\nu},$ (72)
$\mathscr{K}^{(1)\mu}$ $=$ $\frac{1}{p_{n}}\bar{\epsilon}^{\mu\nu\alpha}\bar{p}_{\nu}\mathscr{M}_{\alpha}^{(1)}+\frac{1}{2p_{n}}\Delta^{\mu\lambda}\partial_{\lambda}^{x}{\mathscr{F}}^{(0)}+\frac{1}{2p_{n}}\bar{\mathscr{C}}^{(0)\mu},$ (73)
${\mathscr{P}}^{(1)}$ $=$ $\frac{1}{4p_{n}}\bar{\epsilon}_{\nu\alpha\beta}\partial^{\nu}_{x}{\mathscr{S}}^{(0)\alpha\beta}+\frac{1}{4p_{n}}\bar{\epsilon}_{\nu\alpha\beta}\mathscr{C}^{(0)\nu\alpha\beta}.$ (74)
Substituting Eq.(73) into Eq.(33), we can obtain the antisymmetric Wigner function
$\mathscr{S}^{(1)\mu\nu}=\frac{1}{p_{n}}\epsilon^{\mu\nu\alpha\beta}p_{\alpha}\mathscr{M}_{\beta}^{(1)}+\frac{1}{2p_{n}}\left(\Delta^{\mu\lambda}n^{\nu}-\Delta^{\nu\lambda}n^{\mu}\right)\left(\partial_{\lambda}^{x}{\mathscr{F}}^{(0)}+\mathscr{C}^{(0)}_{\lambda}\right).$ (75)
The difference from the zeroth-order results is that the collision functions have been involved at first order. The Eq.(65) gives the longitudinal constraint condition for ${\mathscr{M}}^{(1)}_{\mu}$
$\bar{p}_{\mu}{\mathscr{M}}^{(1)\mu}$ $=$ $m\sum_{s}s{\mathscr{J}}_{s,n}^{(1)}-\frac{1}{2}\mathscr{C}^{(0)}_{5,n}.$ (76)
Just like at zeroth order, we can decompose $\mathscr{M}^{(1)}_{\mu}$ into the longitudinal and transverse parts $\mathscr{M}^{(1)}_{\mu}=\mathscr{M}^{(1)}_{\parallel\mu}+\mathscr{M}^{(1)}_{\perp\mu}$ with respect to momentum $\bar{p}^{\mu}$ and only transverse component $\mathscr{M}^{\mu}_{\perp}$ is independent.
The Eqs.(59) and (66) together with Eq.(76) give the modification to the on-shell conditions of ${\mathscr{J}}_{s,n}^{(1)}$ and ${\mathscr{M}}^{(1)}_{\mu}$, respectively,
$\left(p^{2}-m^{2}\right)\frac{{\mathscr{J}}_{s,n}^{(1)}}{p_{n}}$ $=$ $-\frac{sm}{4p_{n}}\mathscr{C}^{(0)}_{5,n}-\frac{s}{2p_{n}}\bar{\epsilon}_{\mu\alpha\beta}\bar{p}^{\mu}\mathscr{C}^{(0)\alpha\beta}_{s}-\frac{m}{4p_{n}}n^{\mu}{\mathscr{C}^{(0)\nu}}_{\mu\nu},$ (77)
$\left(p^{2}-m^{2}\right)\frac{{\mathscr{M}}_{\mu}^{(1)}}{p_{n}}$ $=$ $\frac{1}{2}\bar{\mathscr{C}}^{(0)}_{5,\mu}-\frac{\bar{p}_{\mu}}{2p_{n}}{\mathscr{C}}^{(0)}_{5,n}-\frac{1}{2p_{n}}\bar{\epsilon}_{\mu\nu\alpha}\bar{p}^{\nu}\bar{\mathscr{C}}^{(0)\alpha}-\frac{m}{2p_{n}}\bar{\epsilon}_{\mu\alpha\beta}\sum_{s}\mathscr{C}^{(0)\alpha\beta}_{s}.$ (78)
At zeroth order, the remaining Wigner equations are all satisfied automatically. At first order, these equations will result in kinetic equations or constraints on the collision terms. For example, the Eqs.(60) and (70) give the quantum kinetic equations for the zeroth order Wigner functions ${\mathscr{J}}_{s,n}^{(0)}$ and ${\mathscr{M}}^{(0)}_{\mu}$, respectively,
$p^{\mu}\partial^{x}_{\mu}\left(\frac{\mathscr{J}_{s,n}^{(0)}}{p_{n}}\right)$ $=$ $-\frac{ms}{2p_{n}}p^{\nu}\left(\partial^{x}_{\nu}n_{\mu}\right)\frac{{\mathscr{M}}^{(0)\mu}}{p_{n}}-\mathscr{C}^{(0)\mu}_{s,\mu}-\frac{ms}{4p_{n}}\bar{\epsilon}_{\nu\alpha\beta}\mathscr{C}^{(0)\nu\alpha\beta},$ (79)
$p^{\nu}\partial^{x}_{\nu}\left(\frac{\mathscr{M}^{(0)\mu}}{p_{n}}\right)$ $=$ $-\frac{1}{p_{n}}p^{\mu}p^{\nu}\left(\partial^{x}_{\nu}n_{\lambda}\right)\frac{\mathscr{M}^{(0)\lambda}}{p_{n}}-\frac{1}{2}\left(\epsilon^{\mu\nu\alpha\beta}+\frac{p^{\mu}}{p_{n}}\bar{\epsilon}^{\nu\alpha\beta}\right)\mathscr{C}^{(0)}_{\nu\alpha\beta}.$ (80)
Substituting Eqs.(71) and (73) into Eq.(62) and doing some vector algebra together with the kinetic equation (79), we obtain the constraint equation
$\Delta_{\alpha}^{\lambda}\left(\frac{m}{2}{\mathscr{C}}^{(0)\alpha}+\frac{sm}{4}\epsilon^{\alpha\beta\rho\sigma}\mathscr{C}^{(0)}_{\beta\rho\sigma}\right)$ $=$ $\Delta_{\alpha}^{\lambda}\left[p_{\beta}\left(\mathscr{C}^{(0)\alpha\beta}_{s}-\mathscr{C}^{(0)\beta\alpha}_{s}\right)+p^{\alpha}\mathscr{C}^{(0)\beta}_{s,\beta}\right].$ (81)
From the definitions (22-29), we notice that all these collision functions cannot depend on the auxiliary time-like vector $n^{\mu}$. Hence the requirement that the constraint (82) hold for any $n^{\mu}$ lead to the following constraint condition
$\frac{m}{2}{\mathscr{C}}^{(0)\alpha}+\frac{sm}{4}\epsilon^{\alpha\beta\rho\sigma}\mathscr{C}^{(0)}_{\beta\rho\sigma}$ $=$ $p_{\beta}\left(\mathscr{C}^{(0)\alpha\beta}_{s}-\mathscr{C}^{(0)\beta\alpha}_{s}\right)+p^{\alpha}\mathscr{C}^{(0)\beta}_{s,\beta}.$ (82)
Substituting Eqs.(71) and (72) into Eq.(64) and using the kinetic equation (80) and the result (52) give rise to the constraint equation
$mn^{\nu}\epsilon_{\mu\nu\alpha\beta}\sum_{s}s\mathscr{C}^{(0)\alpha\beta}_{s}=n^{\nu}\left(-p_{\mu}{\mathscr{C}^{(0)\lambda}}_{\nu\lambda}+p_{\nu}{\mathscr{C}^{(0)\lambda}}_{\mu\lambda}-\frac{1}{2}\epsilon_{\mu\nu\alpha\beta}p^{\alpha}\epsilon^{\beta\lambda\rho\sigma}{\mathscr{C}^{(0)}}_{\lambda\rho\sigma}\right).$ (83)
The fact that this constraint holds for any $n^{\mu}$ lead to more general constraint condition
$m\epsilon_{\mu\nu\alpha\beta}\sum_{s}s\mathscr{C}^{(0)\alpha\beta}_{s}=-p_{\mu}{\mathscr{C}^{(0)\lambda}}_{\nu\lambda}+p_{\nu}{\mathscr{C}^{(0)\lambda}}_{\mu\lambda}-\frac{1}{2}\epsilon_{\mu\nu\alpha\beta}p^{\alpha}\epsilon^{\beta\lambda\rho\sigma}{\mathscr{C}^{(0)}}_{\lambda\rho\sigma}.$ (84)
Acting the operator $p^{\nu}\partial_{\nu}^{x}$ on the Eq. (53) and using Eqs.(79) and (80) lead to
$2mn^{\mu}p_{\mu}\sum_{s}s\mathscr{C}^{(0)\nu}_{s,\nu}$ $=$ $n^{\mu}\left[p_{\mu}p^{\lambda}\epsilon_{\lambda\nu\alpha\beta}-\left(p^{2}-m^{2}\right)\epsilon_{\mu\nu\alpha\beta}\right]\mathscr{C}^{(0)\nu\alpha\beta}.$ (85)
For the same reason as Eqs.(83), this will lead to more general constraint
$2mp_{\mu}\sum_{s}s\mathscr{C}^{(0)\nu}_{s,\nu}$ $=$ $\left[p_{\mu}p^{\lambda}\epsilon_{\lambda\nu\alpha\beta}-\left(p^{2}-m^{2}\right)\epsilon_{\mu\nu\alpha\beta}\right]\mathscr{C}^{(0)\nu\alpha\beta}.$ (86)
Plugging Eq.(73) into Eq.(67) together with Eqs.(49) and (79) lead to
$m\sum_{s}\mathscr{C}^{(0)\mu}_{s,\mu}$ $=$ $p^{\mu}\mathscr{C}^{(0)}_{\mu}.$ (87)
Multiplying Eq.(76) by $(p^{2}-m^{2})$ and using the Eqs.(77) and (78) gives
$p^{\mu}\mathscr{C}_{5,\mu}^{(0)}$ $=$ $0$ (88)
Acting the operator $p^{\nu}\partial_{\nu}^{x}$ on (55) and multiplying (79) by $(p^{2}-m^{2})$ gives rise to
00 0 $=$ $\left(p^{2}-m^{2}\right)\left(p_{\mu}\mathscr{C}^{(0)\nu}_{s,\nu}-\frac{sm}{4}\epsilon_{\mu\nu\alpha\beta}\mathscr{C}^{(0)\nu\alpha\beta}\right).$ (89)
Making similar manipulation on Eqs.(56) and (80) gives rise to
00 0 $=$ $\left(p^{2}-m^{2}\right)\left(p_{\lambda}\epsilon_{\mu\nu\alpha\beta}-p_{\mu}\epsilon_{\lambda\nu\alpha\beta}\right)\mathscr{C}^{(0)\nu\alpha\beta}.$ (90)
To obtain the final results in Eqs.(89) and (90), we have used the property of the arbitrariness of $n^{\mu}$. However, it is easy to verify that the constraint equation (90) can be derived from Eq.(86) while (89) can be derived from Eqs. (87), (86) and (84). Hence the final independent constraint equations for the collision functions are given by
$p^{\mu}\mathscr{C}_{5,\mu}^{(0)}$ $=$ $0,$ (91)
$m\sum_{s}\mathscr{C}^{(0)\mu}_{s,\mu}$ $=$ $p^{\mu}\mathscr{C}^{(0)}_{\mu},$ (92)
$2mp_{\mu}\sum_{s}s\mathscr{C}^{(0)\nu}_{s,\nu}$ $=$ $\left[p_{\mu}p^{\lambda}\epsilon_{\lambda\nu\alpha\beta}-\left(p^{2}-m^{2}\right)\epsilon_{\mu\nu\alpha\beta}\right]\mathscr{C}^{(0)\nu\alpha\beta},$ (93)
$m\epsilon_{\mu\nu\alpha\beta}\sum_{s}s\mathscr{C}^{(0)\alpha\beta}_{s}$ $=$ $p_{\nu}{\mathscr{C}^{(0)\lambda}}_{\mu\lambda}-p_{\mu}{\mathscr{C}^{(0)\lambda}}_{\nu\lambda}-\frac{1}{2}\epsilon_{\mu\nu\alpha\beta}p^{\alpha}\epsilon^{\beta\lambda\rho\sigma}{\mathscr{C}^{(0)}}_{\lambda\rho\sigma},\hskip 28.45274pt$ (94)
$\frac{m}{2}{\mathscr{C}}^{(0)\alpha}+\frac{sm}{4}\epsilon^{\alpha\beta\rho\sigma}\mathscr{C}^{(0)}_{\beta\rho\sigma}$ $=$ $p_{\beta}\left(\mathscr{C}^{(0)\alpha\beta}_{s}-\mathscr{C}^{(0)\beta\alpha}_{s}\right)+p^{\alpha}\mathscr{C}^{(0)\beta}_{s,\beta}.$ (95)
At first order, we have obtained the constraints for the collision functions and the kinetic equation for zeroth-order independent Wigner functions ${\mathscr{J}}_{s,n}^{(0)}$ and ${\mathscr{M}}^{(0)}_{\mu}$. Since these constraint equations are derived only from the consistency of the Wigner equations, we call them self-consistent constraint equations.
3.3 Second-order result
In order to obtain the kinetic equation for the first-order independent Wigner functions ${\mathscr{J}}_{s,n}^{(1)}$ and ${\mathscr{M}}^{(1)}_{\mu}$, we need to consider the Wigner equations at second order. We can write second-order Wigner equations in the similar form as first order,
$m{\mathscr{F}}^{(2)}$ $=$ $2\left(p_{n}{\mathscr{J}}^{(2)}_{s,n}+\bar{p}_{\mu}\bar{\mathscr{J}}^{(2)\mu}_{s}\right)+\Delta\mathscr{C}_{s,\mu}^{(0)\mu},$ (96)
$-sm{\mathscr{P}}^{(2)}$ $=$ $\partial^{x}_{\mu}{\mathscr{J}}^{(1)\mu}_{s}+\tilde{\mathscr{C}}^{(1)\mu}_{s,\mu},$ (97)
$m{\mathscr{M}}^{(2)}_{\mu}$ $=$ $2s\left(\bar{p}_{\mu}{\mathscr{J}}_{s,n}^{(2)}-p_{n}\bar{\mathscr{J}}_{s,\mu}^{(2)}\right)+\bar{\epsilon}_{\mu\alpha\beta}\partial_{x}^{\alpha}{\mathscr{J}}^{(1)\beta}_{s}+\bar{\epsilon}_{\mu\alpha\beta}\tilde{\mathscr{C}}^{(1)\alpha\beta}_{s},$ (98)
$m\bar{\epsilon}_{\mu\nu\alpha}{\mathscr{K}}^{(2)\alpha}$ $=$ $2s\left(\bar{p}_{\mu}\bar{\mathscr{J}}_{s,\nu}^{(2)}-\bar{p}_{\nu}\bar{\mathscr{J}}_{s,\mu}^{(2)}\right)+\bar{\epsilon}_{\mu\nu\alpha}n_{\beta}\left(\partial_{x}^{\alpha}{\mathscr{J}}^{(1)\beta}_{s}-\partial_{x}^{\beta}{\mathscr{J}}^{(1)\alpha}_{s}\right)$ (99)
$+\bar{\epsilon}_{\mu\nu\alpha}n_{\beta}\left(\tilde{\mathscr{C}}^{(1)\alpha\beta}_{s}-\tilde{\mathscr{C}}^{(1)\beta\alpha}_{s}\right),$
$2m\sum_{s}{\mathscr{J}}_{s,n}^{(2)}$ $=$ $2p_{n}{\mathscr{F}}^{(2)}+n^{\mu}\partial_{x}^{\nu}{\mathscr{S}}_{\mu\nu}^{(1)}+n^{\mu}{\tilde{\mathscr{C}}^{(1)\nu}}_{\ \ \ \ \ \mu\nu},$ (100)
$2m\sum_{s}\bar{\mathscr{J}}_{s,\mu}^{(2)}$ $=$ $2\bar{p}_{\mu}{\mathscr{F}}^{(2)}+\Delta_{\mu}^{\lambda}\partial_{x}^{\nu}{\mathscr{S}}_{\lambda\nu}^{(1)}+\Delta_{\mu}^{\lambda}{\tilde{\mathscr{C}}^{(1)\nu}}_{\ \ \ \ \ \lambda\nu},$ (101)
$-2m\sum_{s}s{\mathscr{J}}_{s,n}^{(2)}$ $=$ $-2\bar{p}_{\mu}{\mathscr{M}}^{(2)\mu}-n^{\mu}\partial^{x}_{\mu}{\mathscr{P}}^{(1)}-\tilde{\mathscr{C}}^{(1)}_{5,n},$ (102)
$-2m\sum_{s}s\bar{\mathscr{J}}_{s,\mu}^{(2)}$ $=$ $2\left(p_{n}\mathscr{M}^{(2)}_{\mu}+\bar{\epsilon}_{\mu\nu\alpha}\bar{p}^{\nu}\mathscr{K}^{(2)\alpha}\right)-\Delta_{\mu}^{\lambda}\partial^{x}_{\lambda}{\mathscr{P}}^{(1)}-\Delta^{\lambda}_{\mu}\tilde{\mathscr{C}}^{(1)}_{5,\lambda},$ (103)
00 0 $=$ $2p^{\nu}{\mathscr{K}}_{\nu}^{(2)}+n^{\mu}\partial_{\mu}^{x}{\mathscr{F}}^{(1)}+\tilde{\mathscr{C}}^{(1)}_{n},$ (104)
00 0 $=$ $-2\left(p_{n}\mathscr{K}^{(2)\mu}-\bar{\epsilon}^{\mu\nu\alpha}\bar{p}_{\nu}\mathscr{M}_{\alpha}^{(2)}\right)+\Delta_{\mu}^{\lambda}\partial_{\lambda}^{x}{\mathscr{F}}^{(1)}+\Delta_{\mu}^{\lambda}\tilde{\mathscr{C}}^{(1)}_{\lambda},$ (105)
00 0 $=$ $2p_{n}{\mathscr{P}}^{(2)}-\frac{1}{2}\bar{\epsilon}_{\nu\alpha\beta}\partial^{\nu}_{x}{\mathscr{S}}^{(1)\alpha\beta}-\frac{1}{2}\bar{\epsilon}_{\nu\alpha\beta}\tilde{\mathscr{C}}^{(1)\nu\alpha\beta},$ (106)
00 0 $=$ $2\bar{p}_{\mu}{\mathscr{P}}^{(2)}+\frac{1}{2}\Delta_{\mu}^{\lambda}\epsilon_{\lambda\nu\alpha\beta}\partial^{\nu}_{x}{\mathscr{S}}^{(1)\alpha\beta}+\frac{1}{2}\Delta_{\mu}^{\lambda}\epsilon_{\lambda\nu\alpha\beta}\tilde{\mathscr{C}}^{(1)\nu\alpha\beta},$ (107)
where we have defined
$\tilde{\mathscr{C}}^{(1)}_{\lambda}$$\equiv$ $\mathscr{C}^{(1)}_{\lambda}-\Delta{\mathscr{C}^{(0)\nu}}_{\lambda\nu},$ (108)
$\tilde{\mathscr{C}}^{(1)}_{5,\lambda}$$\equiv$ $\mathscr{C}^{(1)}_{5,\lambda}-\frac{1}{2}\epsilon_{\lambda\nu\alpha\beta}\Delta\mathscr{C}^{(0)\nu\alpha\beta},$ (109)
$\tilde{\mathscr{C}}^{(1)\alpha\beta}_{s}$$\equiv$ $\mathscr{C}^{(1)\alpha\beta}_{s}-\frac{s}{2}\epsilon^{\alpha\beta\rho\sigma}\Delta\mathscr{C}^{(0)}_{s,\rho\sigma},$ (110)
$\tilde{\mathscr{C}}^{(1)\nu\alpha\beta}$$\equiv$ $\mathscr{C}^{(1)\nu\alpha\beta}+\frac{1}{3}\left(g^{\nu\beta}\Delta\mathscr{C}^{(0)\alpha}-g^{\nu\alpha}\Delta\mathscr{C}^{(0)\beta}\right)+\frac{1}{3}\epsilon^{\nu\alpha\beta\rho}\Delta\mathscr{C}^{(0)}_{5,\rho},$ (111)
because these new collision functions with tilde always appear as a whole.
The Eqs.(98), (100),(105) and (106) expresses other second-order Wigner functions in terms of ${\mathscr{J}}_{s,n}^{(2)}$, ${\mathscr{M}}^{(2)}_{\mu}$,
$\bar{\mathscr{J}}_{s,\mu}^{(2)}$ $=$ $\frac{\bar{p}_{\mu}}{p_{n}}{\mathscr{J}}_{s,n}^{(2)}-\frac{sm}{2p_{n}}{\mathscr{M}}^{(2)}_{\mu}+\frac{s}{2p_{n}}\bar{\epsilon}_{\mu\alpha\beta}\partial_{x}^{\alpha}{\mathscr{J}}^{(1)\beta}_{s}+\frac{s}{2p_{n}}\bar{\epsilon}_{\mu\alpha\beta}\tilde{\mathscr{C}}^{(1)\alpha\beta}_{s},$ (112)
${\mathscr{F}}^{(2)}$ $=$ $\frac{m}{p_{n}}\sum_{s}{\mathscr{J}}_{s,n}^{(2)}-\frac{1}{2p_{n}}n^{\mu}\partial_{x}^{\nu}{\mathscr{S}}_{\mu\nu}^{(1)}-\frac{1}{2p_{n}}n^{\mu}{\tilde{\mathscr{C}}^{(1)\nu}}_{\ \ \ \ \ \mu\nu},$ (113)
$\mathscr{K}^{(2)\mu}$ $=$ $\frac{1}{p_{n}}\bar{\epsilon}^{\mu\nu\alpha}\bar{p}_{\nu}\mathscr{M}_{\alpha}^{(2)}+\frac{1}{2p_{n}}\Delta^{\mu\lambda}\partial_{\lambda}^{x}{\mathscr{F}}^{(1)}+\frac{1}{2p_{n}}\Delta^{\mu\lambda}\tilde{\mathscr{C}}^{(1)}_{\lambda},$ (114)
${\mathscr{P}}^{(2)}$ $=$ $\frac{1}{4p_{n}}\bar{\epsilon}_{\nu\alpha\beta}\partial^{\nu}_{x}{\mathscr{S}}^{(1)\alpha\beta}+\frac{1}{4p_{n}}\bar{\epsilon}_{\nu\alpha\beta}\tilde{\mathscr{C}}^{(1)\nu\alpha\beta}.$ (115)
The antisymmetric tensor Wigner function follows as
$\mathscr{S}^{(2)}_{\mu\nu}$ $=$ $\frac{1}{p_{n}}\epsilon_{\mu\nu\alpha\beta}p^{\alpha}\mathscr{M}^{(2)\beta}+\frac{1}{2p_{n}}\left(\Delta_{\mu}^{\lambda}n_{\nu}-\Delta_{\nu}^{\lambda}n_{\mu}\right)\left(\partial_{\lambda}^{x}{\mathscr{F}}^{(1)}+\tilde{\mathscr{C}}^{(1)}_{\lambda}\right).$ (116)
Substituting the expression (115) into Eqs.(97) and (107), we obtain the kinetic equation for ${\mathscr{J}}_{s,n}^{(1)}$ and ${\mathscr{M}}^{(1)}_{\mu}$, respectively,
$p^{\mu}\partial^{x}_{\mu}\left(\frac{\mathscr{J}_{s,n}^{(1)}}{p_{n}}\right)$ $=$ $-\frac{s}{2p_{n}^{2}}(\partial^{x}_{\nu}n_{\mu})\left[p^{\nu}\left(m{\mathscr{M}}^{(1)\mu}-\bar{\epsilon}^{\mu\rho\sigma}\partial^{x}_{\rho}\mathscr{J}_{s,\sigma}^{(0)}\right)+\bar{\epsilon}^{\mu\nu\beta}p^{\lambda}\left(\mathscr{C}^{(0)}_{s,\beta\lambda}-\mathscr{C}^{(0)}_{s,\lambda\beta}\right)\right]$ (117)
$-\tilde{\mathscr{C}}^{(1)\mu}_{s\mu}-\frac{ms}{4p_{n}}\bar{\epsilon}_{\nu\alpha\beta}\tilde{\mathscr{C}}^{(1)\nu\alpha\beta}{-}\frac{s}{2}\partial^{\mu}_{x}\left(\frac{1}{p_{n}}\bar{\epsilon}_{\mu\alpha\beta}\mathscr{C}^{(0)\alpha\beta}_{s}\right),$
$p^{\nu}\partial_{\nu}^{x}\left(\frac{\mathscr{M}^{(1)\mu}}{p_{n}}\right)$ $=$ $-\frac{1}{p_{n}^{2}}(\partial^{x}_{\lambda}n_{\nu})p^{\lambda}\left(p^{\mu}\mathscr{M}^{(1)\nu}-\frac{1}{2}\bar{\epsilon}^{\mu\nu\rho}\partial^{x}_{\rho}\mathscr{F}^{(0)}-\frac{1}{2}\bar{\epsilon}^{\mu\nu\lambda}p^{\sigma}\mathscr{C}^{(0)}_{\sigma}\right)$ (118)
$-\frac{1}{2}\left(\epsilon^{\mu\nu\alpha\beta}+\frac{p^{\mu}}{p_{n}}\bar{\epsilon}^{\nu\alpha\beta}\right)\left[\tilde{\mathscr{C}}^{(1)}_{\nu\alpha\beta}+\partial^{x}_{\nu}\left(\frac{\bar{\mathscr{C}}^{(0)}_{\alpha}n_{\beta}-\bar{\mathscr{C}}^{(0)}_{\beta}n_{\alpha}}{2p_{n}}\right)\right].$
The equation (102) gives longitudinal constraint condition for ${\mathscr{M}}^{(2)}_{\mu}$
$\bar{p}_{\mu}{\mathscr{M}}^{(2)\mu}$ $=$ $m\sum_{s}s{\mathscr{J}}_{s,n}^{(2)}-\frac{1}{2}\tilde{\mathscr{C}}^{(1)}_{5,n}-\frac{1}{2}n^{\mu}\partial^{x}_{\mu}{\mathscr{P}}^{(1)}.$ (119)
From Eq.(96) and Eq.(103), we obtain the modification to the on-shell conditions of ${\mathscr{J}}_{s,n}^{(2)}$ and ${\mathscr{M}}^{(2)}_{\mu}$, respectively,
$\left(p^{2}-m^{2}\right)\frac{{\mathscr{J}}_{s,n}^{(2)}}{p_{n}}$ (120)
$=$ $-\frac{m}{8p_{n}}n^{\mu}\partial_{x}^{\nu}\left[\frac{\Delta_{\mu}^{\lambda}n_{\nu}-\Delta_{\nu}^{\lambda}n_{\mu}}{p_{n}}\left(\partial_{\lambda}^{x}{\mathscr{F}}^{(0)}+\mathscr{C}^{(0)}_{\lambda}\right)\right]-\frac{m}{4p_{n}}n^{\mu}{\tilde{\mathscr{C}}^{(1)\nu}}_{\ \ \ \ \ \mu\nu}$
$-\frac{sm}{4p_{n}}\left(\tilde{\mathscr{C}}^{(1)}_{5,n}+n^{\mu}\partial^{x}_{\mu}{\mathscr{P}}^{(1)}\right)-\frac{1}{4p_{n}}\bar{\epsilon}_{\mu\alpha\beta}p^{\mu}\partial_{x}^{\alpha}\left[\frac{\bar{\epsilon}^{\beta\rho\sigma}}{p_{n}}\left(\partial^{x}_{\rho}{\mathscr{J}}^{(0)}_{s,\sigma}+\mathscr{C}^{(0)}_{s,\rho\sigma}\right)\right]$
$-\frac{s}{2p_{n}}\bar{\epsilon}_{\mu\alpha\beta}p^{\mu}\tilde{\mathscr{C}}^{(1)\alpha\beta}_{s}-\frac{1}{2}\Delta\mathscr{C}_{s,\mu}^{(0)\mu},$
$\left(p^{2}-m^{2}\right)\frac{\mathscr{M}^{(2)}_{\mu}}{p_{n}}$ (121)
$=$ $\frac{1}{2}\Delta_{\mu}^{\lambda}\partial^{x}_{\lambda}{\mathscr{P}}^{(1)}+\frac{1}{2}\Delta^{\lambda}_{\mu}\tilde{\mathscr{C}}^{(1)}_{5,\lambda}-\frac{m}{4p_{n}}\sum_{s}\bar{\epsilon}_{\mu\alpha\beta}\partial_{x}^{\alpha}\left[\frac{s\bar{\epsilon}^{\beta\rho\sigma}}{p_{n}}\left(\partial^{x}_{\rho}{\mathscr{J}}^{(0)}_{s,\sigma}+\mathscr{C}^{(0)}_{s,\rho\sigma}\right)\right]$
$-\frac{m}{2p_{n}}\bar{\epsilon}_{\mu\alpha\beta}\sum_{s}\tilde{\mathscr{C}}^{(1)\alpha\beta}_{s}-\frac{\bar{p}_{\mu}}{2p_{n}}\left(\tilde{\mathscr{C}}^{(1)}_{5,n}+n^{\nu}\partial^{x}_{\nu}{\mathscr{P}}^{(1)}\right)$
$+\frac{1}{4p_{n}}\bar{\epsilon}_{\mu\nu\alpha}\bar{p}^{\nu}\Delta^{\alpha\lambda}\partial_{\lambda}^{x}\left[\frac{n^{\mu}}{p_{n}}\left(\partial_{x}^{\nu}{\mathscr{S}}_{\mu\nu}^{(0)}+{\mathscr{C}^{(0)\nu}}_{\mu\nu}\right)\right]-\frac{1}{2p_{n}}\bar{\epsilon}_{\mu\nu\alpha}\bar{p}^{\nu}\tilde{\mathscr{C}}^{(1)\alpha}.$
Following the same procedure as we have taken at first order, we can obtain the self-consistent constraints for the collision terms at the second order. Besides much more complicated vector algebraic and derivative operation at second order, we also encounter the terms associated with the derivative terms with $n^{\mu}$ such as $\partial_{x}^{\nu}n^{\mu}$ in the second-order constraint equations. However, it is remarkable that all these terms cancel each other if we impose the first-order constraint equations (91)-(95). The final independent constraint equations are given by
$p^{\mu}\tilde{\mathscr{C}}^{(1)}_{5,\mu}$ $=$ $\frac{1}{4}\epsilon^{\mu\nu\alpha\beta}\partial_{\mu}^{x}\mathscr{C}^{(0)}_{\nu\alpha\beta}-m\sum_{s}s\Delta\mathscr{C}_{s,\mu}^{(0)\mu},$ (122)
$m\sum_{s}\tilde{\mathscr{C}}^{(1)\lambda}_{s,\lambda}$ $=$ $p^{\lambda}\tilde{\mathscr{C}}^{(1)}_{\lambda}-\frac{1}{2}\partial_{x}^{\nu}{\mathscr{C}^{(0)\lambda}}_{\nu\lambda},$ (123)
$2mp_{\mu}\sum_{s}s\tilde{\mathscr{C}}^{(1)\lambda}_{s,\lambda}$ $=$ $[p_{\mu}p^{\lambda}\epsilon_{\lambda\nu\alpha\beta}-(p^{2}-m^{2})\epsilon_{\mu\nu\alpha\beta}]\tilde{\mathscr{C}}^{(1)\nu\alpha\beta}$ (124)
$+\epsilon_{\mu\lambda\alpha\beta}\partial^{\lambda}_{x}(m\sum_{s}\mathscr{C}^{(0)\alpha\beta}_{s}-p^{\beta}\mathscr{C}^{(0)\alpha})-p^{\nu}\partial_{\nu}^{x}\mathscr{C}^{(0)}_{5,\mu},$
$m\epsilon_{\mu\nu\alpha\beta}\sum_{s}s\tilde{\mathscr{C}}^{(1)\alpha\beta}_{s}$ $=$ $p_{\nu}{\tilde{\mathscr{C}}^{(1)\lambda}}_{\ \ \ \ \ \mu\lambda}-p_{\mu}{\tilde{\mathscr{C}}^{(1)\lambda}}_{\ \ \ \ \ \nu\lambda}-\frac{1}{2}\epsilon_{\mu\nu\rho\sigma}p^{\rho}\epsilon^{\sigma\lambda\alpha\beta}\tilde{\mathscr{C}}^{(1)}_{\lambda\alpha\beta}$ (125)
$+\frac{1}{2}(\partial^{x}_{\nu}\mathscr{C}^{(0)}_{\mu}-\partial^{x}_{\mu}\mathscr{C}^{(0)}_{\nu})-\frac{1}{2}\epsilon_{\mu\nu\rho\sigma}\partial_{x}^{\rho}{\mathscr{C}}^{(0)\sigma}_{5},$
$\frac{m}{2}\tilde{\mathscr{C}}^{(1)\mu}+\frac{sm}{4}\epsilon^{\mu\beta\rho\sigma}\tilde{\mathscr{C}}^{(1)}_{\beta\rho\sigma}$ $=$ $p_{\lambda}(\tilde{\mathscr{C}}^{(1)\mu\lambda}_{s}-\tilde{\mathscr{C}}^{(1)\lambda\mu}_{s})+p^{\mu}\tilde{\mathscr{C}}^{(1)\lambda}_{s,\lambda}-\frac{s}{2}\epsilon^{\mu\beta\rho\sigma}\partial_{\beta}^{x}\mathscr{C}^{(0)}_{s,\rho\sigma}.$ (126)
4 Self-consistent solutions for collision terms
As we all know, these collision terms cannot be determined by a group of closed equations due to BBGKY hierarchy. Some proper approximation must be imposed to make the equations closed. No matter what approximation we make, the self-consistent constraints derived in previous sections should be fulfilled. In this section, we will find self-consistent specific solutions for these collision terms.
Let us start from the first-order constraints conditions (91)-(95) and try to find a specific expression as simple as possible but still not very trivial. In the simplest case, we will assume that all the collision terms at the first order are on-shell. It is easy to verify that the following expressions always satisfy all first-order constraints conditions (91)-(95),
$\mathscr{C}_{5,\mu}^{(0)}$ $=$ $0,$ (127)
$\mathscr{C}^{(0)\mu}$ $=$ $m\sum_{s}\mathscr{X}_{s}^{(0)\mu},$ (128)
$\mathscr{C}^{(0)\mu\nu}_{s}$ $=$ $\mathscr{X}_{s}^{(0)\mu}p^{\nu},$ (129)
$\mathscr{C}^{(0)\nu\alpha\beta}$ $=$ $-\frac{m}{3}\epsilon^{\mu\nu\alpha\beta}\sum_{s}s\mathscr{X}^{(0)}_{s,\mu},$ (130)
where $\mathscr{X}_{s}^{(0)\mu}$ with chirality index $s$ can be arbitrary vector function with the on-shell condition,
$\left(p^{2}-m^{2}\right){\mathscr{X}}_{s}^{(0)}=0,\ \ \ \textrm{or equivalently, }\ \ \ {\mathscr{X}}_{s}^{(0)}={\mathcal{X}}_{s}^{(0)}\delta(p^{2}-m^{2}).$ (131)
The physical meaning of function $\mathscr{X}_{s}^{(0)\mu}$ depends on the specific system we are considering. In the chiral limit $m=0$, we find that only $\mathscr{C}^{(0)\mu\nu}_{s}$ survive and $\mathscr{C}^{(0)\mu}$ and $\mathscr{C}^{(0)\nu\alpha\beta}$ both vanish and are trivial. In order to make $\mathscr{C}^{(0)\nu\alpha\beta}$ not trivial in the chiral limit because it is related to the magnetic moment distribution, we can find a slightly more complex solution
$\mathscr{C}_{5\mu}^{(0)}$ $=$ $0,$ (132)
$\mathscr{C}^{(0)\mu}$ $=$ $m\sum_{s}\mathscr{X}_{s}^{(0)\mu},$ (133)
$\mathscr{C}^{(0)\mu\nu}_{s}$ $=$ $\mathscr{X}_{s}^{(0)\mu}p^{\nu}+\frac{sm}{4u\cdot p}\left(u^{\mu}\mathscr{Y}^{(0)\nu}-u^{\nu}\mathscr{Y}^{(0)\mu}\right),$ (134)
$\mathscr{C}^{(0)\nu\alpha\beta}$ $=$ $-\frac{u^{\nu}}{u\cdot p}\epsilon^{\alpha\beta\rho\sigma}p_{\rho}\mathscr{Y}^{(0)}_{\sigma}-\frac{m}{3}\epsilon^{\mu\nu\alpha\beta}\sum_{s}s\mathscr{X}^{(0)}_{s,\mu},$ (135)
where we have introduced the vector functions $u^{\mu}$ and $\mathscr{Y}^{(0)\mu}$ which satisfy
$u^{2}=1,\ \ \ u\cdot\mathscr{Y}^{(0)}=0,\ \ \ p\cdot\mathscr{Y}^{(0)}=0,$ (136)
and the on-shell condition
$\left(p^{2}-m^{2}\right){\mathscr{Y}}^{(0)}=0,\ \ \ \textrm{or equivalently, }\ \ \ {\mathscr{Y}}^{(0)}={\mathcal{Y}}^{(0)}\delta(p^{2}-m^{2}).$ (137)
In the chiral limit, we have
$\mathscr{C}_{5,\mu}^{(0)}$ $=$ $0,$ (138)
$\mathscr{C}^{(0)\mu}$ $=$ $0,$ (139)
$\mathscr{C}^{(0)\mu\nu}_{s}$ $=$ $\mathscr{X}_{s}^{(0)\mu}p^{\nu},$ (140)
$\mathscr{C}^{(0)\nu\alpha\beta}$ $=$ $-\frac{1}{u\cdot p}u^{\nu}\epsilon^{\alpha\beta\rho\sigma}p_{\rho}\mathscr{Y}^{(0)}_{\sigma}.$ (141)
It should be pointed out that the normalized time-like vector $u^{\mu}$ should be regarded as a physical function which could depend on both coordinate and momentum and has no relations to $n^{\mu}$. With this specific expressions, we find that these collision terms do not modify either longitudinal constraint condition for ${\mathscr{M}}^{(1)}_{\mu}$
$\bar{p}_{\mu}{\mathscr{M}}^{(1)\mu}$ $=$ $m\sum_{s}s{\mathscr{J}}_{s,n}^{(1)},$ (142)
or the on-shell conditions of ${\mathscr{J}}_{s,n}^{(1)}$ and ${\mathscr{M}}^{(1)}_{\mu}$
$\left(p^{2}-m^{2}\right)\frac{{\mathscr{J}}_{s,n}^{(1)}}{p_{n}}$ $=$ $0,$ (143)
$\left(p^{2}-m^{2}\right)\frac{{\mathscr{M}}_{\mu}^{(1)}}{p_{n}}$ $=$ $0$ (144)
The general expressions are given by
${\mathscr{J}}_{s,n}^{(1)}$ $=$ $p_{n}{\mathcal{J}}_{s,n}^{(1)}\delta(p^{2}-m^{2}),$ (145)
${\mathscr{M}}_{\mu}^{(1)}$ $=$ $p_{n}{\mathcal{M}}_{\mu}^{(1)}\delta(p^{2}-m^{2}).$ (146)
The quantum kinetic equations of ${\mathscr{J}}_{s,n}^{(0)}$ and ${\mathscr{M}}^{(0)}_{\mu}$ with these collision terms follows
$p^{\mu}\partial^{x}_{\mu}\left(\frac{\mathscr{J}_{sn}^{(0)}}{p_{n}}\right)$ $=$ $-\frac{ms}{2p_{n}}p^{\nu}\left(\partial^{x}_{\nu}n_{\mu}\right)\frac{{\mathscr{M}}^{(0)\mu}}{p_{n}}$ (147)
$-\mathscr{X}^{(0)\mu}_{s}p_{\mu}+\frac{sm^{2}}{2p_{n}}\sum_{s^{\prime}}s^{\prime}\mathscr{X}^{(0)n}_{s^{\prime}}-\frac{ms}{2p_{n}}\mathscr{Y}_{n}^{(0)},$
$p^{\nu}\partial^{x}_{\nu}\left(\frac{\mathscr{M}^{(0)\mu}}{p_{n}}\right)$ $=$ $-\frac{1}{p_{n}}p^{\mu}p^{\nu}\left(\partial^{x}_{\nu}n_{\lambda}\right)\frac{\mathscr{M}^{(0)\lambda}}{p_{n}}$ (148)
$-\frac{1}{p_{n}}\left(p^{\mu}\mathscr{Y}_{n}^{(0)}-p_{n}\mathscr{Y}^{(0)\mu}\right)-m\sum_{s^{\prime}}s^{\prime}\left(\mathscr{X}^{(0)\mu}_{s^{\prime}}-\frac{p^{\mu}}{p_{n}}\mathscr{X}^{(0)n}_{s^{\prime}}\right).$
From the second-order constraints (122)-(126), we can find a second-order solution
$\tilde{\mathscr{C}}_{5,\mu}^{(1)}$ $=$ $\frac{u_{\mu}}{2u\cdot p}\left(-\partial^{\nu}_{x}\mathscr{Y}^{(0)}_{\nu}+m\sum_{s}s\partial^{\nu}_{x}\mathscr{X}^{(0)}_{s,\nu}\right),$ (149)
$\tilde{\mathscr{C}}^{(1)\mu}$ $=$ $m\sum_{s}\mathscr{X}_{s}^{(1)\mu}-\epsilon^{\mu\nu\alpha\beta}\partial^{x}_{\nu}\left(\frac{u_{\alpha}\mathscr{Y}_{\beta}^{(0)}}{2u\cdot p}\right),$ (150)
$\tilde{\mathscr{C}}^{(1)\mu\nu}_{s}$ $=$ $\mathscr{X}_{s}^{(1)\mu}p^{\nu}+\frac{sm}{4u\cdot p}\left(u^{\mu}\mathscr{Y}^{(1)\nu}-u^{\nu}\mathscr{Y}^{(1)\mu}\right),$ (151)
$\tilde{\mathscr{C}}^{(1)\nu\alpha\beta}$ $=$ $-\frac{1}{u\cdot p}u^{\nu}\epsilon^{\alpha\beta\rho\sigma}p_{\rho}\mathscr{Y}^{(1)}_{\sigma}-\frac{m}{3}\epsilon^{\mu\nu\alpha\beta}\sum_{s}s\mathscr{X}^{(1)}_{s,\mu}.$ (152)
With these specific expressions, first-order Wigner functions have the following form
${\mathscr{J}}_{s,\mu}^{(1)}$ $=$ $\frac{p_{\mu}}{p_{n}}{\mathscr{J}}_{s,n}^{(1)}-\frac{sm}{2p_{n}}{\mathscr{M}}^{(1)}_{\mu}+\frac{s}{2p_{n}}\bar{\epsilon}_{\mu\alpha\beta}\partial_{x}^{\alpha}{\mathscr{J}}^{(0)\beta}_{s}$ (153)
$+\frac{1}{2p_{n}}\bar{\epsilon}_{\mu\alpha\beta}\left(s\mathscr{X}_{s}^{(0)\alpha}p^{\beta}+\frac{m}{2u\cdot p}u^{\alpha}\mathscr{Y}^{(0)\beta}\right),$
${\mathscr{F}}^{(1)}$ $=$ $\frac{m}{p_{n}}\sum_{s}{\mathscr{J}}_{s,n}^{(1)}-\frac{1}{2p_{n}}n^{\mu}\partial_{x}^{\nu}{\mathscr{S}}_{\mu\nu}^{(0)}-\frac{1}{2p_{n}(u\cdot p)}u_{\nu}\bar{\epsilon}^{\nu\rho\sigma}p_{\rho}\mathscr{Y}^{(0)}_{\sigma},$ (154)
${\mathscr{P}}^{(1)}$ $=$ $\frac{1}{4p_{n}}\bar{\epsilon}_{\nu\alpha\beta}\partial^{\nu}_{x}{\mathscr{S}}^{(0)\alpha\beta}+\frac{1}{2p_{n}}\left(\mathscr{Y}^{(0)}_{n}-m\sum_{s}s\mathscr{X}_{sn}^{(0)}\right),$ (155)
$\mathscr{K}^{(1)\mu}$ $=$ $\frac{1}{p_{n}}\bar{\epsilon}^{\mu\nu\alpha}\bar{p}_{\nu}\mathscr{M}_{\alpha}^{(1)}+\frac{1}{2p_{n}}\Delta^{\mu\lambda}\partial_{\lambda}^{x}{\mathscr{F}}^{(0)}+\frac{m}{2p_{n}}\sum_{s}\bar{\mathscr{X}}_{s}^{(0)\mu}.$ (156)
The quantum kinetic equations of ${\mathscr{J}}_{s,n}^{(1)}$ and ${\mathscr{M}}^{(1)}_{\mu}$ are given by
$p^{\mu}\partial^{x}_{\mu}\left(\frac{\mathscr{J}_{s,n}^{(1)}}{p_{n}}\right)$ $=$ $-\frac{s}{2p_{n}}p^{\nu}(\partial^{x}_{\nu}n_{\mu})\left(m\frac{{\mathscr{M}}^{(1)\mu}}{p_{n}}-\frac{1}{p_{n}}\bar{\epsilon}^{\mu\alpha\beta}\partial^{x}_{\alpha}\mathscr{J}_{s,\beta}^{(0)}\right)$ (157)
$+\frac{s}{2p_{n}^{2}}p^{\nu}(\partial^{x}_{\nu}n_{\mu})\bar{\epsilon}^{\mu\alpha\beta}p_{\beta}\mathscr{X}_{s,\alpha}^{(0)}-\frac{s}{2p_{n}}\bar{\epsilon}^{\mu\alpha\beta}p_{\beta}\partial^{x}_{\mu}\mathscr{X}_{s,\alpha}^{(0)}$
$-\left(\mathscr{X}^{(1)\mu}_{s}p_{\mu}-\frac{sm^{2}}{2p_{n}}\sum_{s^{\prime}}s^{\prime}\mathscr{X}^{(1)n}_{s^{\prime}}\right)-\frac{ms}{2p_{n}}\mathscr{Y}_{n}^{(1)},$
$p^{\nu}\partial^{x}_{\nu}\left(\frac{\mathscr{M}^{(1)\mu}}{p_{n}}\right)$ $=$ $-\frac{1}{p_{n}}p^{\lambda}(\partial_{\lambda}^{x}n_{\nu})\left(p^{\mu}\frac{\mathscr{M}^{(1)\nu}}{p_{n}}-\frac{1}{2p_{n}}\bar{\epsilon}^{\mu\nu\alpha}\partial_{\alpha}^{x}{\mathscr{F}}^{(0)}\right)$ (158)
$+\frac{m}{2p_{n}^{2}}p^{\lambda}(\partial_{\lambda}^{x}n_{\nu})\bar{\epsilon}^{\mu\nu\alpha}\sum_{s^{\prime}}\mathscr{X}^{(0)}_{s^{\prime},\alpha}-\frac{m}{2p_{n}}\bar{\epsilon}^{\mu\nu\alpha}\partial_{\nu}^{x}\sum_{s^{\prime}}\mathscr{X}^{(0)}_{s^{\prime},\alpha}$
$-m\sum_{s^{\prime}}s^{\prime}\left(\mathscr{X}^{(1)\mu}_{s^{\prime}}-\frac{p^{\mu}}{p_{n}}\mathscr{X}^{(1)n}_{s^{\prime}}\right)-\frac{1}{p_{n}}\left(p^{\mu}\mathscr{Y}_{n}^{(1)}-p_{n}\mathscr{Y}^{(1)\mu}\right).$
5 Self-consistent relaxation-time approximation
The general kinetic equation is very difficult to tackle because the BBGKY hierarchy or the non-linear collision terms. The relaxation-time approximation has been proposed since 1950s Bhatnagar:1954zz ; Marle:1969 ; Anderson:1974nyl and used quite successfully in several field physics. Recently, the quantum kinetic equation at naive relaxation-time approximation has been discussed in Hidaka:2017auj ; Wang:2021qnt . In this section, we will present the quantum kinetic equation at self-consistent relaxation-time approximation, which is consistent with the self-consistent constraints obtained in previous sections.
From the requirement (136), we can assume the collision terms $\mathscr{X}_{s\mu}^{(k)}$ and $\mathscr{Y}_{\mu}^{(k)}$ ($k=0,1$) take the conventional form at relaxation-time approximation
$\mathscr{X}_{s\mu}^{(k)}=\frac{1}{p_{u}}\left(\frac{u_{\mu}}{\tau_{1s}}+\frac{p_{\mu}}{p_{u}\tau_{2s}}\right)\delta{\mathscr{J}}_{s,u}^{(k)}+\frac{ms}{2p_{u}^{2}\tau_{3}}\delta{\mathscr{M}}_{u\perp,\mu}^{(k)},\ \ \ \mathscr{Y}_{\mu}^{(k)}=-\frac{1}{\tau_{4}}\delta{\mathscr{M}}_{u\perp,\mu}^{(k)},$ (159)
where $\tau_{1s}$, $\tau_{2s}$ , $\tau_{3}$, and $\tau_{4}$ denote the relaxation time parameters associtated with distribution functions ${\mathscr{J}}_{s,n}$ and ${\mathscr{M}}_{\perp,\mu}$. In this section, we identify the time-like vector $u^{\mu}$ as the fluid velocity. We have decomposed the functions along $u^{\mu}$ instead of $n^{\mu}$ because all the collision functions cannot depends on $n^{\mu}$ and they can only depend on some physical quantity such as $u^{\mu}$. The symbols in the above expressions are defined by
$p_{u}=u\cdot p,\ \ \ \ {\mathscr{J}}_{s,u}^{(k)}=u\cdot{\mathscr{J}}_{s}^{(k)},\ \ \ {\mathscr{M}}_{u\perp,\mu}^{(k)}=\frac{1}{2}\epsilon_{\mu\nu\alpha\beta}u^{\nu}\mathscr{S}^{(k)\alpha\beta},$ (160)
$\delta\mathscr{J}_{s,u}^{(k)}={\mathscr{J}}_{s,u}^{(k)}-{\mathscr{J}}_{s,u,\textrm{eq}}^{(k)},\ \ \ \delta{\mathscr{M}}_{u\perp,\mu}^{(k)}={\mathscr{M}}_{u\perp,\mu}^{(k)}-{\mathscr{M}}_{u\perp,\textrm{eq},\mu}^{(k)},$ (161)
where the subscript ‘eq’ indicates the corresponding functions at local or global equilibrium. With these expressions, the quantum kinetic equations are given by
$p^{\mu}\partial^{x}_{\mu}\left(\frac{\mathscr{J}_{sn}^{(k)}}{p_{n}}\right)$ $=$ $-\frac{s}{2p_{n}}p^{\nu}\left(\partial^{x}_{\nu}n_{\mu}\right)\left(m\frac{{\mathscr{M}}^{(k)\mu}}{p_{n}}-\frac{1}{p_{n}}\bar{\epsilon}^{\mu\alpha\beta}\partial^{x}_{\alpha}\mathscr{J}_{s\beta}^{(k-1)}\right)$ (162)
$-\frac{p_{\mu}}{p_{u}}\left(\frac{u^{\mu}}{\tau_{1s}}+\frac{p^{\mu}}{p_{u}\tau_{2s}}\right)\delta{\mathscr{J}}_{s,u}^{(k)}+\frac{sm^{2}}{2p_{n}p_{u}}\sum_{s^{\prime}}s^{\prime}\left(\frac{n\cdot u}{\tau_{1s^{\prime}}}+\frac{p_{n}}{p_{u}\tau_{2s^{\prime}}}\right)\delta{\mathscr{J}}_{s^{\prime},u}^{(k)}$
$+\frac{s}{2p_{n}^{2}}\bar{\epsilon}^{\mu\alpha\beta}p_{\beta}\left[p^{\nu}(\partial^{x}_{\nu}n_{\mu})-p_{n}\partial^{x}_{\mu}\right]\left(\frac{u_{\alpha}}{p_{u}\tau_{1s}}\delta{\mathscr{J}}_{s,u}^{(k-1)}\right)$
$+\frac{ms}{2p_{n}p_{u}}\left(\frac{p_{u}}{\tau_{4}}+\frac{m^{2}}{p_{u}\tau_{3}}\right)n^{\mu}\delta{\mathscr{M}}_{u\perp,\mu}^{(k)}$
$+\frac{m}{4p_{n}^{2}}\bar{\epsilon}^{\mu\alpha\beta}p_{\beta}\left[p^{\nu}(\partial^{x}_{\nu}n_{\mu})-p_{n}\partial^{x}_{\mu}\right]\left(\frac{1}{p_{u}^{2}\tau_{3}}{\mathscr{M}}_{u\perp,\alpha}^{(k-1)}\right),$
$p^{\nu}\partial^{x}_{\nu}\left(\frac{\mathscr{M}^{(k)\mu}}{p_{n}}\right)$ $=$ $-\frac{1}{p_{n}}p^{\lambda}(\partial_{\lambda}^{x}n_{\nu})\left(p^{\mu}\frac{\mathscr{M}^{(k)\nu}}{p_{n}}-\frac{1}{2p_{n}}\bar{\epsilon}^{\mu\nu\alpha}\partial_{\alpha}^{x}{\mathscr{F}}^{(k-1)}\right)$ (163)
$+\frac{p^{\mu}n_{\lambda}-p_{n}g^{\mu}_{\lambda}}{p_{n}p_{u}}\left(\frac{p_{u}}{\tau_{4}}+\frac{m^{2}}{p_{u}\tau_{3}}\right)\delta{\mathscr{M}}_{u\perp}^{(k)\lambda}$
$+\frac{m\left(p^{\mu}n_{\lambda}-p_{n}g^{\mu}_{\lambda}\right)}{p_{n}p_{u}}\sum_{s^{\prime}}s^{\prime}\left(\frac{u^{\lambda}}{\tau_{1s^{\prime}}}+\frac{p^{\lambda}}{p_{u}\tau_{2s^{\prime}}}\right)\delta{\mathscr{J}}_{s^{\prime},u}^{(k)}$
$+\frac{m}{2p_{n}^{2}}\bar{\epsilon}^{\mu\nu\alpha}\left[p^{\lambda}(\partial_{\lambda}^{x}n_{\nu})-p_{n}\partial_{\nu}^{x}\right]\sum_{s^{\prime}}\frac{1}{p_{u}}\left(\frac{u_{\alpha}}{\tau_{1s^{\prime}}}+\frac{p_{\alpha}}{p_{u}\tau_{2s^{\prime}}}\right)\delta{\mathscr{J}}_{s^{\prime},u}^{(k-1)}.$
It should be noted that when $k=0$ we define the functions with superscript $k=-1$ to vanish. The zeroth-order Wigner functions include no collision contribution. The Wigner functions at first order read
${\mathscr{J}}_{s,\mu}^{(1)}$ $=$ $\frac{p_{\mu}}{p_{n}}{\mathscr{J}}_{s,n}^{(1)}-\frac{sm}{2p_{n}}{\mathscr{M}}^{(1)}_{\mu}+\frac{s}{2p_{n}}\bar{\epsilon}_{\mu\alpha\beta}\partial_{x}^{\alpha}{\mathscr{J}}^{(0)\beta}_{s}$ (164)
$+\frac{s}{2p_{n}p_{u}\tau_{1s}}\bar{\epsilon}_{\mu\alpha\beta}u^{\alpha}p^{\beta}\delta{\mathscr{J}}_{s^{\prime},u}^{(0)}+\frac{m}{4p_{n}p_{u}^{2}}\bar{\epsilon}_{\mu\alpha\beta}\left(\frac{p^{\beta}}{\tau_{3}}+\frac{p_{u}u^{\beta}}{\tau_{4}}\right)\delta{\mathscr{M}}_{u\perp}^{(0)\alpha},$
${\mathscr{F}}^{(1)}$ $=$ $\frac{m}{p_{n}}\sum_{s}{\mathscr{J}}_{s,n}^{(1)}-\frac{1}{2p_{n}}n^{\mu}\partial_{x}^{\nu}{\mathscr{S}}_{\mu\nu}^{(0)}+\frac{1}{2p_{n}p_{u}\tau_{4}}u_{\nu}\bar{\epsilon}^{\nu\rho\sigma}p_{\rho}\delta{\mathscr{M}}_{u\perp,\sigma}^{(0)},$ (165)
${\mathscr{P}}^{(1)}$ $=$ $\frac{1}{4p_{n}}\bar{\epsilon}_{\nu\alpha\beta}\partial^{\nu}_{x}{\mathscr{S}}^{(0)\alpha\beta}-\frac{m}{2p_{n}p_{u}}\sum_{s}s\left(\frac{n\cdot u}{\tau_{1s}}+\frac{p_{n}}{p_{u}\tau_{2s}}\right)\delta{\mathscr{J}}_{s,u}^{(0)}$ (166)
$-\frac{1}{2p_{n}}\left(\frac{1}{\tau_{4}}+\frac{m^{2}}{p_{u}^{2}\tau_{3}}\right)n^{\mu}\delta{\mathscr{M}}_{u\perp,\mu}^{(0)},$
$\mathscr{S}^{(1)\mu\nu}$ $=$ $\frac{1}{p_{n}}\epsilon^{\mu\nu\alpha\beta}p_{\alpha}\mathscr{M}_{\beta}^{(1)}+\frac{1}{2p_{n}}\left(\Delta^{\mu\lambda}n^{\nu}-\Delta^{\nu\lambda}n^{\mu}\right)\partial_{\lambda}^{x}{\mathscr{F}}^{(0)}$ (167)
$+\frac{m}{2p_{n}p_{u}}\left(\Delta^{\mu\lambda}n^{\nu}-\Delta^{\nu\lambda}n^{\mu}\right)\sum_{s}\left(\frac{u_{\lambda}}{\tau_{1s}}+\frac{p_{\lambda}}{p_{u}\tau_{2s}}\right)\delta{\mathscr{J}}_{s,u}^{(0)},$
${\mathscr{M}}^{(1)\mu}_{\parallel}$ $=$ $\frac{m\bar{p}^{\mu}}{\bar{p}^{2}}\sum_{s}s{\mathscr{J}}_{s,n}^{(1)}.$ (168)
These kinetic equations and Wigner functions can be simplified if we make the arbitrary subsidiary $n^{\mu}$ and the physical fluid velocity $u^{\mu}$ coincide. For brevity, we will replace $u^{\mu}$ with $n^{\mu}$ instead of replacing $n^{\mu}$ with $u^{\mu}$. Besides, we also sum the zeroth and first orders into a unified form by defining
$\mathscr{J}_{sn}=\mathscr{J}_{sn}^{(0)}+\mathscr{J}_{sn}^{(1)},\ \ \ \mathscr{M}^{\mu}=\mathscr{M}^{(0)\mu}+\mathscr{M}^{(1)\mu}.$ (169)
With $n^{\mu}=u^{\mu}$, the quantum kinetic equations given above in separate orders are equivalent to the following form up to the first order $\hbar$,
$p^{\mu}\partial^{x}_{\mu}\left(\frac{\mathscr{J}_{sn}}{p_{n}}\right)$ $=$ $-\frac{s}{2p_{n}}p^{\nu}\left(\partial^{x}_{\nu}n_{\mu}\right)\left[m\frac{{\mathscr{M}}^{\mu}}{p_{n}}-\frac{1}{p_{n}}\bar{\epsilon}^{\mu\alpha\beta}\partial^{x}_{\alpha}\left(\frac{p_{\beta}}{p_{n}}{\mathscr{J}}_{s,n}-\frac{sm}{2p_{n}}{\mathscr{M}}_{\beta}\right)\right]$ (170)
$-\left(\frac{1}{\tau_{1s}}+\frac{m^{2}}{p_{n}^{2}\tau_{2s}}\right)\delta{\mathscr{J}}_{s,n}+\frac{sm^{2}}{2p_{n}^{2}}\sum_{s^{\prime}}s^{\prime}\left(\frac{1}{\tau_{1s^{\prime}}}+\frac{1}{\tau_{2s^{\prime}}}\right)\delta{\mathscr{J}}_{s^{\prime},n}$
$-\frac{s}{2p_{n}^{2}\tau_{1s}}\bar{\epsilon}^{\mu\alpha\beta}p_{\beta}(\partial^{x}_{\mu}n_{\alpha})\delta{\mathscr{J}}_{s,n}$
$+\frac{m}{4p_{n}^{2}}\bar{\epsilon}^{\mu\alpha\beta}p_{\beta}\left[p^{\nu}(\partial^{x}_{\nu}n_{\mu})-p_{n}\partial^{x}_{\mu}\right]\left(\frac{\delta{\mathscr{M}}_{\perp\alpha}}{p_{n}^{2}\tau_{3}}\right),$
$p^{\nu}\partial^{x}_{\nu}\left(\frac{\mathscr{M}^{\mu}}{p_{n}}\right)$ $=$ $-\frac{1}{p_{n}}p^{\lambda}(\partial_{\lambda}^{x}n_{\nu})\left[p^{\mu}\frac{\mathscr{M}^{\nu}}{p_{n}}-\frac{m}{2p_{n}}\bar{\epsilon}^{\mu\nu\alpha}\partial_{\alpha}^{x}\left(\frac{1}{p_{n}}\sum_{s}{\mathscr{J}}_{s,n}\right)\right]$ (171)
$-\left(\frac{1}{\tau_{4}}+\frac{m^{2}}{p_{n}^{2}\tau_{3}}\right)\delta{\mathscr{M}}^{\mu}_{\perp}+\frac{m\bar{p}^{\mu}}{p_{n}^{2}}\sum_{s^{\prime}}\frac{s^{\prime}}{\tau_{1s^{\prime}}}\delta{\mathscr{J}}_{s^{\prime},n}$
$+\frac{m}{2p_{n}^{2}}\bar{\epsilon}^{\mu\nu\alpha}\left[p^{\lambda}(\partial_{\lambda}^{x}n_{\nu})-p_{n}\partial_{\nu}^{x}\right]\sum_{s^{\prime}}\frac{p_{\alpha}}{p_{n}^{2}\tau_{2s^{\prime}}}\delta{\mathscr{J}}_{s^{\prime},n}.$
The Wigner functions are given by
${\mathscr{J}}_{s,\mu}$ $=$ $\frac{p_{\mu}}{p_{n}}{\mathscr{J}}_{s,n}-\frac{sm}{2p_{n}}{\mathscr{M}}_{\mu}+\frac{s}{2p_{n}}\bar{\epsilon}_{\mu\alpha\beta}p^{\beta}\partial_{x}^{\alpha}\left(\frac{{\mathscr{J}}_{s,n}}{p_{n}}\right)+\frac{m}{4p_{n}^{3}\tau_{3}}\bar{\epsilon}_{\mu\alpha\beta}p^{\beta}\delta{\mathscr{M}}^{\alpha}_{\perp},$ (172)
${\mathscr{F}}$ $=$ $\frac{m}{p_{n}}\sum_{s}{\mathscr{J}}_{s,n}-\frac{1}{2p_{n}}\epsilon_{\mu\nu\alpha\beta}p^{\alpha}n^{\mu}\partial_{x}^{\nu}\left(\frac{{\mathscr{M}}^{\beta}}{p_{n}}\right),$ (173)
${\mathscr{P}}$ $=$ $-\frac{1}{2p_{n}}\left(p_{\nu}n_{\sigma}-p_{n}g_{\nu\sigma}\right)\partial^{\nu}_{x}\left(\frac{{\mathscr{M}}^{\sigma}}{p_{n}}\right)-\frac{m}{2p_{n}^{2}}\sum_{s}s\left(\frac{1}{\tau_{1s}}+\frac{1}{\tau_{2s}}\right)\delta{\mathscr{J}}_{s,n},$ (174)
$\mathscr{S}^{\mu\nu}$ $=$ $\frac{1}{p_{n}}\epsilon^{\mu\nu\alpha\beta}p_{\alpha}\mathscr{M}_{\beta}+\frac{m}{2p_{n}}\left(\Delta^{\mu\lambda}n^{\nu}-\Delta^{\nu\lambda}n^{\mu}\right)\partial_{\lambda}^{x}\sum_{s}\frac{{\mathscr{J}}_{s,n}}{p_{n}}$ (175)
$+\frac{m}{2p_{n}^{3}}\left(\bar{p}^{\mu}n^{\nu}-\bar{p}^{\nu}n^{\mu}\right)\sum_{s}\frac{1}{\tau_{2s}}\delta{\mathscr{J}}_{s,n},$
${\mathscr{M}}^{\mu}_{\parallel}$ $=$ $\frac{m\bar{p}^{\mu}}{\bar{p}^{2}}\sum_{s}s{\mathscr{J}}_{s,n}.$ (176)
The Wigner functions have been greatly reduced when we identify $n^{\mu}$ as the local fluid velocity $u^{\mu}$. The scalar Wigner function ${\mathscr{F}}$ even has no explicit dependence on the collision terms. From (181) and (9), we can obtain the axial and vector Wigner functions directly
${\mathscr{A}}_{\mu}$ $=$ $\frac{p_{\mu}}{p_{n}}{\mathscr{A}}_{n}-\frac{m}{p_{n}}{\mathscr{M}}_{\mu}+\frac{1}{2p_{n}}\bar{\epsilon}_{\mu\alpha\beta}p^{\beta}\partial_{x}^{\alpha}\left(\frac{{\mathscr{V}}_{n}}{p_{n}}\right),$ (177)
${\mathscr{V}}_{\mu}$ $=$ $\frac{p_{\mu}}{p_{n}}{\mathscr{V}}_{n}+\frac{1}{2p_{n}}\bar{\epsilon}_{\mu\alpha\beta}p^{\beta}\partial_{x}^{\alpha}\left(\frac{{\mathscr{A}}_{n}}{p_{n}}\right)+\frac{m}{2p_{n}^{3}\tau_{3}}\bar{\epsilon}_{\mu\alpha\beta}p^{\beta}\delta{\mathscr{M}}^{\alpha}_{\perp}$ (178)
which denote the spin polarization distribution and electric charge separation in phase space, respectively. We notice that the spin polarization represented by ${\mathscr{A}}_{\mu}$ has no explicit dependence on the collision terms in local comoving frame. The collision dependence has already been totally coded in the distribution function itself. The electric charge separation represented by ${\mathscr{V}}_{\mu}$ has dependence on the collisions terms only from the the transverse magnetic momentum distribution $\delta{\mathscr{M}}^{\alpha}_{\perp}$. This contribution is totally from non-equilibrium effect and could affect chiral magnetic effect when there exists transverse polarization in relativistic heavy-ion collisions. Certainly, these conclusions are drawn from relaxation-time approximation and might be changed when we consider more general collision contribution. However it still shed light on how the collision terms contribute to the spin polarization or electric charge separation in heavy-ion collisions.
In the chiral limit, we can safely set $m=0$ in our formalism of GCKT and obtain the quantum kinetic equations
$p^{\mu}\partial^{x}_{\mu}\left(\frac{\mathscr{J}_{s,n}}{p_{n}}\right)$ $=$ $\frac{s}{2p_{n}^{2}}p^{\nu}\left(\partial^{x}_{\nu}n_{\mu}\right)\bar{\epsilon}^{\mu\alpha\beta}p_{\beta}\partial^{x}_{\alpha}\left(\frac{\mathscr{J}_{s,n}}{p_{n}}\right)$ (179)
$-\frac{1}{\tau_{1s}}\left[1+\frac{s}{2p_{n}^{2}}\bar{\epsilon}^{\mu\alpha\beta}p_{\beta}(\partial^{x}_{\mu}n_{\alpha})\right]\delta{\mathscr{J}}_{s,n},$
$p^{\nu}\partial^{x}_{\nu}\left(\frac{\mathscr{M}^{\mu}_{\perp}}{p_{n}}\right)$ $=$ $-\frac{1}{p_{n}}p^{\lambda}(\partial_{\lambda}^{x}n_{\nu})p^{\mu}\frac{\mathscr{M}^{\nu}_{\perp}}{p_{n}}-\frac{1}{\tau_{4}}\delta{\mathscr{M}}^{\mu}_{\perp}$ (180)
and the Wigner functions read
${\mathscr{J}}_{s,\mu}$ $=$ $\frac{p_{\mu}}{p_{n}}{\mathscr{J}}_{s,n}+\frac{s}{2p_{n}}\bar{\epsilon}_{\mu\alpha\beta}p^{\beta}\partial_{x}^{\alpha}\left(\frac{{\mathscr{J}}_{s,n}}{p_{n}}\right),$ (181)
${\mathscr{F}}$ $=$ $-\frac{1}{2p_{n}}\epsilon_{\mu\nu\alpha\beta}p^{\alpha}n^{\mu}\partial_{x}^{\nu}\left(\frac{{\mathscr{M}}^{\beta}}{p_{n}}\right),$ (182)
${\mathscr{P}}$ $=$ $-\frac{1}{2p_{n}}\left(p_{\nu}n_{\sigma}-p_{n}g_{\nu\sigma}\right)\partial^{\nu}_{x}\left(\frac{{\mathscr{M}}^{\sigma}}{p_{n}}\right),$ (183)
$\mathscr{S}^{\mu\nu}$ $=$ $\frac{1}{p_{n}}\epsilon^{\mu\nu\alpha\beta}p_{\alpha}\mathscr{M}_{\beta},$ (184)
$\mathscr{M}^{\mu}_{\parallel}$ $=$ 00 0 (185)
We see that the Wigner functions $\mathscr{J}_{s,n}$ and $\mathscr{M}^{\mu}$ are completely decoupled with each other as it should be. It is remarkable that all the Wigner functions have no explicit dependence on the collision terms at chiral limit.
6 Summary and discussion
In this paper, we start from the gauge-invariant Wigner function formalism in quantum electrodynamics and derive the quantum kinetic theory with BBGKY hierarchy or general collisions terms. With the help of the semiclassical $\hbar$ expansion, we use the property of self-consistency of the Wigner equations to constrain the collision terms up to second order. These constraint equations provide general necessary conditions to make some approximation or simplification in some specific cases just like we impose the conservation laws in the collision terms. When only three vector functions involved, we present a specific solution for the collision terms in the formalism of GCKT. We further specialize this solution in the form of relaxation-time approximation and present the self-consistent quantum kinetic theory at relaxation-time approximation which is compatible with the self-consistent constraints. We find the quantum kinetic theory at relaxation-time approximation can be greatly simplified by defining or decomposing the distribution functions in the comoving fluid frame with velocity $u^{\mu}$. Some Wigner functions exhibit even no explicit dependence on the collision terms and the implicit dependence is totally coded into the distribution functions when we write the equations or functions in the fluid comoving frame.
Although we didn’t take into account the mean electromagnetic field during deriving the quantum kinetic equation, Wigner functions, and the constraint conditions with collision terms, it is straightforward and easy to obtain the results when a mean field is involved. As we mentioned in the section 2, the collision functions discussed in this work actually contain all possible terms, including mean field contribution. At mean field approximation, all the constraint conditions derived in previous sections will be satisfied automatically as verified in Gao:2018wmr ; Gao:2019znl ; Ma:2022ins . Therefore, we can simply decompose the general electromagnetic field in (5) into mean field part and quantized field part, separate the mean field part from the quantized field, and directly obtain the quantum kinetic equation and Wigner functions with mean electromagnetic field involved. The constraint conditions for collision terms from the quantized field part remain unchanged.
We have presented the quantum kinetic equations and Wigner equations in 8-dimensional phase space, 4-dimensional coordinate space $x^{\mu}$ plus 4-dimensional momentum space $p^{\mu}$. In 8-dimensional phase space, the on-shell Dirac delta function is always involved. We can integrate the time-like component of momentum to eliminate the singular Dirac delta function and obtain the quantum kinetic theory in 7-dimensional phase space, which can be applied to make numerical calculation directly. Actually, it is a trivial task to obtain the 7-dimensional quantum kinetic theory from the 8-dimensional one in sections 4 and 5 because only the onshell Dirac delta functions are involved and there are no derivative terms with respect to the momentum when the background electromagnetic field is absent.
It will be valuable to solve the self-consistent quantum kinetic equation at relaxation-time approximation analytical or numerically. We will postpone these interesting and valuable studies in the future.
7 Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant Nos. 12175123 and 12321005.

### v2

@@
 Abstract
-We derive self-consistent constraint conditions for collision terms in quantum kinetic theory within Wigner function formalism.
-We present simple specific solutions for the collision terms in quantum kinetic theory which is compatible with these constraints.
-We obtain the self-consistent quantum kinetic theory at relaxation-time approximation, with which we discuss spin polarization effects and electric charge separation effects in relativistic heavy-ion collisions.
-1 Introduction
+We derive self-consistent constraint conditions for collision terms in quantum kinetic theory using the Wigner function formalism.
+We present specific solutions for these collision terms that align with the constraints.
+we develop quantum kinetic theory at relaxation-time approximation.
+With this framework, we discuss spin polarization and electric charge separation effects in relativistic heavy-ion collisions.
+I Introduction
 In 1872, Boltzmann, the founder of the kinetic theory, derived his famous classical kinetic equation from the physical point of view Boltzmann:1872 .
@@
 In the 2000s, the pioneering work Liang:2004ph ; Liang:2004xn ; Kharzeev:2007jp ; Fukushima:2008xe ; Kharzeev:2007tn ; Gao:2007bc ; Becattini:2007sr initiated the research on the chiral and spin effects in relativistic heavy-ion collisions, which lead to further development on the derivation of quntum kinetic theory with chiral or spin degree of freedom Gao:2012ix ; Stephanov:2012ki ; Son:2012zy ; Chen:2012ca ; Manuel:2013zaa ; Manuel:2014dza ; Chen:2014cla ; Chen:2015gta ; Hidaka:2016yjf ; Mueller:2017lzw ; Huang:2018wdl ; Gao:2018wmr ; Liu:2018xip ; Gao:2019znl ; Weickgenannt:2019dks ; Hattori:2019ahi ; Wang:2019moi ; Lin:2019fqo ; Li:2019qkf ; Sheng:2020oqs ; Guo:2020zpa ; Weickgenannt:2020aaf ; Hayata:2020sqz ; Weickgenannt:2021cuo ; Sheng:2021kfc ; Fauth:2021nwe ; Luo:2021uog ; ChenShiLe:2021guf ; Lin:2021mvw ; Chen:2021azy ; Sheng:2022ssd ; Fang:2022ttm ; Wagner:2022amr ; Ma:2022ins ; Wagner:2023cct .
-Collision terms in kinetic equation play a central role in describing the evolution from non-equilibrium to equilibrium, and can be derived from the BBGKY equation or hierarchyBogolyubov:1946 by introducing certain additional assumptions.
+Collision terms in kinetic equation play a central role in describing the evolution in non-equilibrium systume, and can be derived from the BBGKY equation or hierarchyBogolyubov:1946 by introducing certain additional assumptions.
 As we all know, in general, the collision terms should satisfy the principle of detailed balance and the constraints of charge conservation or energy-momentum conservation.
@@
 We choose natural units such that $\hbar=c=1$ unless otherwise stated.
-2 Wigner function formalism
+II Wigner function formalism
 The quantum kinetic theory can be built with different formalisms.
@@
 where we have recovered the $\hbar$ dependence so that we can make semiclassical expansion in the next section.
-3 Self-consistent constraints on collision terms
+III Self-consistent constraints on collision terms
 The Wigner equations given in Eqs.(13)-(19) are very complicated and totally coupled with each other.
@@
 $\bar{\epsilon}_{\mu\alpha\beta}=\epsilon_{\mu\nu\alpha\beta}n^{\nu}.$ (35)
-3.1 Zeroth-order result
+III.1 Zeroth-order result
 At zeroth order, all the collision functions vanish in the Wigner equations Eqs.(13)-(19).
@@
 At zeroth order, we have no either constraints for the collision function or the kinetic equation for independent Wigner functions ${\mathscr{J}}_{s,n}^{(0)}$ and ${\mathscr{M}}^{(0)}_{\mu}$.
-3.2 First-order result
+III.2 First-order result
 The collision terms and kinetic equation begin to appear at first order. Similar to the procedure we carried out at zeroth order, the first-order Wigner equations can be obtained from Eqs.(13)-(19)
@@
 $\frac{m}{2}{\mathscr{C}}^{(0)\alpha}+\frac{sm}{4}\epsilon^{\alpha\beta\rho\sigma}\mathscr{C}^{(0)}_{\beta\rho\sigma}$ $=$ $p_{\beta}\left(\mathscr{C}^{(0)\alpha\beta}_{s}-\mathscr{C}^{(0)\beta\alpha}_{s}\right)+p^{\alpha}\mathscr{C}^{(0)\beta}_{s,\beta}.$ (95)
-At first order, we have obtained the constraints for the collision functions and the kinetic equation for zeroth-order independent Wigner functions ${\mathscr{J}}_{s,n}^{(0)}$ and ${\mathscr{M}}^{(0)}_{\mu}$.
+At first order, we have obtained the constraints for the zeroth-order collision functions and the kinetic equation for zeroth-order independent Wigner functions ${\mathscr{J}}_{s,n}^{(0)}$ and ${\mathscr{M}}^{(0)}_{\mu}$.
 Since these constraint equations are derived only from the consistency of the Wigner equations, we call them self-consistent constraint equations.
-3.3 Second-order result
+III.3 Second-order result
 In order to obtain the kinetic equation for the first-order independent Wigner functions ${\mathscr{J}}_{s,n}^{(1)}$ and ${\mathscr{M}}^{(1)}_{\mu}$, we need to consider the Wigner equations at second order.
@@
 $+\frac{1}{4p_{n}}\bar{\epsilon}_{\mu\nu\alpha}\bar{p}^{\nu}\Delta^{\alpha\lambda}\partial_{\lambda}^{x}\left[\frac{n^{\mu}}{p_{n}}\left(\partial_{x}^{\nu}{\mathscr{S}}_{\mu\nu}^{(0)}+{\mathscr{C}^{(0)\nu}}_{\mu\nu}\right)\right]-\frac{1}{2p_{n}}\bar{\epsilon}_{\mu\nu\alpha}\bar{p}^{\nu}\tilde{\mathscr{C}}^{(1)\alpha}.$
-Following the same procedure as we have taken at first order, we can obtain the self-consistent constraints for the collision terms at the second order.
+Following the same procedure as we have taken at first order, we can obtain the self-consistent constraints for the collision terms at the first order.
 Besides much more complicated vector algebraic and derivative operation at second order, we also encounter the terms associated with the derivative terms with $n^{\mu}$ such as $\partial_{x}^{\nu}n^{\mu}$ in the second-order constraint equations.
-However, it is remarkable that all these terms cancel each other if we impose the first-order constraint equations (91)-(95).
+However, it is remarkable that all these terms cancel each other if we impose the zeroth-order constraint equations (91)-(95).
 The final independent constraint equations are given by
@@
 $\frac{m}{2}\tilde{\mathscr{C}}^{(1)\mu}+\frac{sm}{4}\epsilon^{\mu\beta\rho\sigma}\tilde{\mathscr{C}}^{(1)}_{\beta\rho\sigma}$ $=$ $p_{\lambda}(\tilde{\mathscr{C}}^{(1)\mu\lambda}_{s}-\tilde{\mathscr{C}}^{(1)\lambda\mu}_{s})+p^{\mu}\tilde{\mathscr{C}}^{(1)\lambda}_{s,\lambda}-\frac{s}{2}\epsilon^{\mu\beta\rho\sigma}\partial_{\beta}^{x}\mathscr{C}^{(0)}_{s,\rho\sigma}.$ (126)
-4 Self-consistent solutions for collision terms
+IV Some other constraint conditions
+In addition to the constraint equations for zeroth order (91)-(95) and first order (122)-(126) which are derived from the self-consistency of the Wigner equations, we can obtain other constraint equations from the specific structure defined from (22) to (29).
+The antisymmetry of the electromagnetic field tensor $F^{\mu\lambda}$ requires the following constraints
+$\partial_{\nu}^{p}{\mathscr{C}^{\nu}}$ $=$ $0,\ \ \ \partial_{\nu}^{p}{\mathscr{C}^{\nu}_{5}}=0,\ \ \ \partial_{\mu}^{p}\mathscr{C}^{\mu\nu}_{s}=0,\ \ \ \partial_{\nu}^{p}{\mathscr{C}^{\nu\alpha\beta}}=0,$ (127)
+$\partial_{\nu}^{p}\Delta{\mathscr{C}^{\nu}}$ $=$ $0\displaystyle 0,\ \ \partial_{\nu}^{p}\Delta{\mathscr{C}^{\nu}_{5}}=0,\ \ \partial_{\mu}^{p}\Delta\mathscr{C}^{\mu\nu}_{s}=0,\ \ \partial_{\nu}^{p}\Delta{\mathscr{C}^{\nu\alpha\beta}}=$ 0 .
+(128)
+The constraints (127) require that the zeroth-order ${\mathscr{C}^{(0)\nu}}$, ${\mathscr{C}^{(0)\nu}_{5}}$, $\mathscr{C}^{(0)\mu\nu}_{s}$ and ${\mathscr{C}^{(0)\nu\alpha\beta}}$ in (91)-(95) should satisfy the same constraints.
+However, the constraints (127) and (128) cannot lead to the similar constraints on the first-order ${\tilde{\mathscr{C}}^{(1)\nu}}$, ${\tilde{\mathscr{C}}^{(1)\nu}_{5}}$, $\tilde{\mathscr{C}}^{(1)\mu\nu}_{s}$ and $\tilde{\mathscr{C}}^{(1)\nu\alpha\beta}$ in (127) and (128) due to the out-of-step linear combination in the definitions (108)-(111).
+In the following sections, we will disregard these constraints and only focus on the constraints (91)-(95) and (122)-(126).
+V Self-consistent specific solutions for collision terms
 As we all know, these collision terms cannot be determined by a group of closed equations due to BBGKY hierarchy.
 Some proper approximation must be imposed to make the equations closed.
-No matter what approximation we make, the self-consistent constraints derived in previous sections should be fulfilled.
-In this section, we will find self-consistent specific solutions for these collision terms.
-Let us start from the first-order constraints conditions (91)-(95) and try to find a specific expression as simple as possible but still not very trivial.
+No matter what approximation we make, the self-consistent constraints (91)-(95) and (122)-(126) derived in previous sections should be fulfilled.
+In this section, we will find self-consistent particular solutions for these collision terms.
+Let us start from the zeroth-order constraints conditions (91)-(95) and try to find a specific expression as simple as possible but still not very trivial.
 In the simplest case, we will assume that all the collision terms at the first order are on-shell.
@@
 $-\frac{1}{p_{n}}\left(p^{\mu}\mathscr{Y}_{n}^{(0)}-p_{n}\mathscr{Y}^{(0)\mu}\right)-m\sum_{s^{\prime}}s^{\prime}\left(\mathscr{X}^{(0)\mu}_{s^{\prime}}-\frac{p^{\mu}}{p_{n}}\mathscr{X}^{(0)n}_{s^{\prime}}\right).$
-From the second-order constraints (122)-(126), we can find a second-order solution
+From the second-order constraints (122)-(126), we can find a first-order solution
 $\tilde{\mathscr{C}}_{5,\mu}^{(1)}$ $=$ $\frac{u_{\mu}}{2u\cdot p}\left(-\partial^{\nu}_{x}\mathscr{Y}^{(0)}_{\nu}+m\sum_{s}s\partial^{\nu}_{x}\mathscr{X}^{(0)}_{s,\nu}\right),$ (151)
@@
 $-m\sum_{s^{\prime}}s^{\prime}\left(\mathscr{X}^{(1)\mu}_{s^{\prime}}-\frac{p^{\mu}}{p_{n}}\mathscr{X}^{(1)n}_{s^{\prime}}\right)-\frac{1}{p_{n}}\left(p^{\mu}\mathscr{Y}_{n}^{(1)}-p_{n}\mathscr{Y}^{(1)\mu}\right).$
-5 Self-consistent relaxation-time approximation
+VI Self-consistent relaxation-time approximation
 The general kinetic equation is very difficult to tackle because the BBGKY hierarchy or the non-linear collision terms.
@@
 It is remarkable that all the Wigner functions have no explicit dependence on the collision terms at chiral limit.
-6 Summary and discussion
+VII Summary and discussion
 In this paper, we start from the gauge-invariant Wigner function formalism in quantum electrodynamics and derive the quantum kinetic theory with BBGKY hierarchy or general collisions terms.
 With the help of the semiclassical $\hbar$ expansion, we use the property of self-consistency of the Wigner equations to constrain the collision terms up to second order.
-These constraint equations provide general necessary conditions to make some approximation or simplification in some specific cases just like we impose the conservation laws in the collision terms.
+These constraint equations provide general necessary conditions to make some approximation or simplification in some specific cases.
 When only three vector functions involved, we present a specific solution for the collision terms in the formalism of GCKT.
-We further specialize this solution in the form of relaxation-time approximation and present the self-consistent quantum kinetic theory at relaxation-time approximation which is compatible with the self-consistent constraints.
+We further specialize this solution in the form of relaxation-time approximation and present the self-consistent quantum kinetic theory at relaxation-time approximation which aligns with the self-consistent constraints.
 We find the quantum kinetic theory at relaxation-time approximation can be greatly simplified by defining or decomposing the distribution functions in the comoving fluid frame with velocity $u^{\mu}$.
@@
 Although we didn’t take into account the mean electromagnetic field during deriving the quantum kinetic equation, Wigner functions, and the constraint conditions with collision terms, it is straightforward and easy to obtain the results when a mean field is involved.
-As we mentioned in the section 2, the collision functions discussed in this work actually contain all possible terms, including mean field contribution.
+As we mentioned in the section II, the collision functions discussed in this work actually contain all possible terms, including mean field contribution.
 At mean field approximation, all the constraint conditions derived in previous sections will be satisfied automatically as verified in Gao:2018wmr ; Gao:2019znl ; Ma:2022ins .
@@
 We can integrate the time-like component of momentum to eliminate the singular Dirac delta function and obtain the quantum kinetic theory in 7-dimensional phase space, which can be applied to make numerical calculation directly.
-Actually, it is a trivial task to obtain the 7-dimensional quantum kinetic theory from the 8-dimensional one in sections 4 and 5 because only the onshell Dirac delta functions are involved and there are no derivative terms with respect to the momentum when the background electromagnetic field is absent.
+Actually, it is a trivial task to obtain the 7-dimensional quantum kinetic theory from the 8-dimensional one in sections V and VI because only the onshell Dirac delta functions are involved and there are no derivative terms with respect to the momentum when the background electromagnetic field is absent.
 It will be valuable to solve the self-consistent quantum kinetic equation at relaxation-time approximation analytical or numerically.
 We will postpone these interesting and valuable studies in the future.
-7 Acknowledgments
+VIII Acknowledgments
 This work was supported in part by the National Natural Science Foundation of China under Grant Nos. 12175123 and 12321005.